### v1

# InvBench: Can LLMs Accelerate Program Verification with Invariant Synthesis?

Abstract
Program verification relies on loop invariants, yet automatically discovering strong invariants remains a long-standing challenge. We introduce a principled framework for evaluating LLMs on invariant synthesis. Our approach uses a verifier-based decision procedure with a formal soundness guarantee and assesses not only correctness but also the speedup that invariants provide in verification. We evaluate 7 state-of-the-art LLMs, and existing LLM-based verifiers against the traditional solver UAutomizer. While LLM-based verifiers represent a promising direction, they do not yet offer a significant advantage over UAutomizer. Model capability also proves critical, as shown by sharp differences in speedups across models, and our benchmark remains an open challenge for current LLMs. Finally, we show that supervised fine-tuning and Best-of-N sampling can improve performance: fine-tuning on 3589 instances raises the percentage of speedup cases for Qwen3-Coder-480B from 8% to 29.2%, and Best-of-N sampling with N=16 improves Claude-sonnet-4 from 8.8% to 22.1%.
1 Introduction
Program verification aims to provide formal guarantees that software behaves as intended, with applications in many safety-critical domains (Fan et al., 2017; Luckcuck et al., 2019). A long-standing challenge in this area, studied for more than four decades, is the automatic discovery of loop invariants. In this work, we investigate whether large language models (LLMs) can accelerate program verification by generating useful loop invariants.
Loop invariants are conditions that hold before and after each loop iteration, and they are central to deductive program verification. To accelerate program verification, loop invariants must not only be correct but also sufficiently strong to prove the assertions. Generating correct invariants is relatively easy, since any universally true condition qualifies. However, only strong invariants can reduce verification effort and lead to a speedup. For example, in Figure˜1, the invariant $x>0$ is correct but not strong enough to prove the final assertion $x\neq 145$, whereas $x\equiv 3\pmod{7}$ is both correct and sufficiently strong.
Discovering such invariants is difficult and undecidable in general, which has motivated a long line of research. Traditional approaches include constraint solving (Colon et al., 2003; Gupta et al., 2009), dynamic analysis (Le et al., 2019), etc. Since invariant discovery is undecidable in general, researchers have tried a variety of learning-based methods (Li et al., 2017; Ezudheen et al., 2018). Building on this progression, the strong capabilities of LLMs in code generation and program reasoning (Austin et al., 2021; Chen et al., 2021; Wei et al., 2025b; c) naturally motivate a systematic evaluation of their potential for invariant discovery.
Pei et al. (2023) is the first work to evaluate the capabilities of LLMs in invariant generation. However, their methodology considers only correctness and does not assess how strong the generated invariants are. As a result, LLMs may generate correct invariants that perform well under their evaluation metric but provide no benefit for accelerating the verification process in real-world settings. Furthermore, their notion of correctness is not based on formal verification. Instead, it is determined by direct comparison with invariants generated by an existing tool, namely Daikon (Ernst et al., 2007). Daikon is a dynamic analysis tool whose invariants are not guaranteed to be sound, since they are inferred from observed test executions rather than proven across all possible executions. As a result, the ground-truth invariants themselves may be incorrect. Moreover, directly comparing LLM-generated invariants with Daikon’s output can lead to rejecting many correct invariants. For example, if an LLM proposes $a>0$ for an integer $a$ while Daikon reports $a\geq 1$, the evaluation in Pei et al. (2023) would incorrectly classify the LLM’s result as wrong, even though the two are equivalent. Hence, prior work’s evaluation methodology cannot reliably capture the correctness of LLM-generated invariants, let alone how useful they are for verification.
A series of follow-up works inspired by Pei et al. (2023) have proposed LLM-based verifiers (Wu et al., 2024b; a; Kamath et al., 2024). Instead of evaluating LLMs in isolation, these efforts develop verification frameworks powered by LLMs. However, the evaluation of such tools suffers from two key limitations: (1) each work introduces a customized dataset and reports results only on that dataset, making cross-comparison impossible; and (2) many omit direct comparison with state-of-the-art traditional verifiers, underscoring the need for evaluation against established baselines.
Figure 1: Illustration of InvBench’s evaluation pipeline. The LLM proposes an invariant by specifying a program location and predicate (e.g., location B with x % 7 == 3). The verification procedure then incorporates this invariant to prove the property x != 145 using two verifier queries, and we measure the resulting speedup relative to a baseline without LLM assistance.
Our work, InvBench, introduces a principled methodology for evaluating the capabilities of LLMs in invariant synthesis. Instead of checking whether LLMs reproduce invariants discovered by other tools, we employ a verifier-based decision procedure that directly determines the correctness of LLM-generated invariants, and we prove this procedure to be sound. We formalize the methodology as a proof calculus, providing a rigorous foundation for invariant evaluation. Unlike previous work that focuses on optimizing verification performance by designing tailored interactive protocols between solvers and LLMs (Chakraborty et al., 2023; Wu et al., 2024b; Kamath et al., 2024; Wu et al., 2024a), our goal is to develop a simple, first-order verification procedure suitable for evaluating the invariant generation capability of LLMs. Our formalization provides such a framework, offering a general solution for invariant evaluation.
Since the purpose of invariant synthesis is to accelerate verification, invariants that are too weak to aid verification or too difficult to verify as correct offer little practical value. To capture this, we evaluate the invariants by measuring the speedup they provide in the overall verification process.
To support comparison across solvers and LLMs, we construct a dataset of 226 instances derived from the most recent edition of the software verification competition SV-Comp (Beyer & Strejček, 2025) and use it to evaluate multiple LLM-based verifiers on this common benchmark. In addition, we assess the state-of-the-art traditional (i.e., non-LLM-based) verifier UAutomizer (Schüssele et al., 2024) both on our dataset and on each of the customized datasets introduced in prior work (Wu et al., 2024b; Kamath et al., 2024; Wu et al., 2024a).
UAutomizer consistently outperforms prior LLM-based verifiers on both our dataset and their customized datasets across all settings, with the only exception of LEMUR (Wu et al., 2024b), which was specifically designed for problems that UAutomizer fails to solve. These results suggest that while LLM-based verifiers are a promising direction, they do not yet offer a significant advantage over non-LLM-based approaches in general. Our findings further highlight the importance of using established datasets and baselines for meaningful progress in the community. They also highlight the role of the underlying solver, suggesting that base solver choice can influence outcomes.
In addition, we evaluate 7 state-of-the-art LLMs. The results indicate that model capability is critical, as shown by the sharp differences in achieved speedups. Nonetheless, LLMs remain largely unable to solve the hard subset of InvBench, underscoring that our dataset presents an open challenge for current LLMs.
We also investigate the extent to which supervised fine-tuning and Best-of-N sampling can enhance model performance in accelerating verification. To this end, we construct a fine-tuning dataset of 3589 instances and show that training on this dataset raises the percentage of speedup cases for Qwen3-Coder-480B from 8% to 29.2%. Similarly, Best-of-N sampling with N = 16 improves Claude-sonnet-4 from 8.8% to 22.1%.
In summary, our contributions are as follows:
  • We propose a verification procedure for evaluating the invariant generation capabilities of LLMs, assessing both correctness and their effectiveness in accelerating verification.
  • We evaluate 7 state-of-the-art LLMs, and provide comparisons between existing LLM-based verifiers and the state-of-the-art non-LLM-based solver UAutomizer.
  • We construct a dataset of 3589 instances for training. We demonstrate that both supervised fine-tuning and Best-of-N sampling can improve model performance in accelerating verification.
2 Related Work
Traditional Methods for Program Invariant Generation.
A long line of research has explored invariant synthesis using traditional techniques without machine learning, including model checking (Flanagan & Qadeer, 2002; Lahiri & Bryant, 2007; Hojjat & Rümmer, 2018; Vediramana Krishnan et al., 2024), abstract interpretation (Karr, 1976; Cousot & Cousot, 1977; Cousot & Halbwachs, 1978; Cousot & Cousot, 1979), constraint solving (Gulwani et al., 2009; Gupta et al., 2009), Craig interpolation (Jhala & McMillan, 2006; McMillan, 2010), and syntax-guided synthesis (Fedyukovich & Bodík, 2018). Prior work evaluating LLM-generated invariants (Pei et al., 2023) has relied on Daikon (Ernst et al., 2007), a tool for dynamic invariant detection (Echenim et al., 2019; Le et al., 2019). Daikon executes the program, observes runtime values, and reports properties that consistently hold over the observed executions. However, such invariants may fail to generalize to all possible executions, thereby compromising soundness. Our approach instead employs a verifier-based decision procedure relying on UAutomizer (Schüssele et al., 2024) that ensures soundness.
Learning-Based Method for Invariant Generation
Machine learning based techniques have been widely adopted in invariant synthesis, including decision tree (Garg et al., 2014; 2016; Ezudheen et al., 2018; Riley & Fedyukovich, 2022; Xu et al., 2020), support vector machine (Li et al., 2017; Sharma et al., 2012), reinforcement learning (Si et al., 2018; Yu et al., 2023), and others (Sharma et al., 2013; Ryan et al., 2019; Yao et al., 2020). More recently, large language models have demonstrated strong capabilities in reasoning about code and logic (Wei et al., 2025a; b; d), giving rise to a series of work that explore using LLMs for finding invariants. Pei et al. (2023) is the first pioneering work that evaluates LLMs’ capabilities in finding invariants, but it is not a sound evaluation. Chakraborty et al. (2023) proposes ranking the invariants generated by LLMs. LaM4Inv (Wu et al., 2024a) proposes a “query-filter-reassemble” strategy to leverage the generation power of LLMs, LEMUR (Wu et al., 2024b) proposes a sound algorithm with back-tracking. Loopy (Kamath et al., 2023) builds an LLM-based pipeline that combines clause sampling, Houdini filtering, and error-guided repair. Liu et al. (2024) proposes a new dataset including problems from courses and real-world systems. Our work explores fine-tuning and the best-of-N sampling method to improve LLMs’ capabilities.
3 Method
3.1 Preliminary
We formalize the task of loop invariant synthesis using standard Hoare logic Hoare (1969). A Hoare triple $\{P\}\,S\,\{Q\}$ specifies that if the precondition $P$ holds before executing a statement $S$, then the postcondition $Q$ will hold after its execution. In the context of loops, an invariant $I$ is a logical proposition that summarizes the state of the program at each iteration, and it is the key to proving the validity of Hoare triples involving loops. For a loop of the form while $B$ do $S$, the goal of invariant synthesis is to identify a loop invariant $I$ that satisfies the following inference rule:
$\frac{P\Rightarrow I\quad\{I\wedge B\}\,S\,\{I\}\quad I\wedge\lnot B\Rightarrow Q}{\{P\}\ \texttt{while}\ B\ \texttt{do}\ S\ \{Q\}}$
Here, $P$ is the precondition, $Q$ is the postcondition, $B$ is the loop condition, and $S$ is the loop body. Intuitively, the inference rule requires that the invariant $I$ holds at the beginning of the loop ($P\Rightarrow I$), is preserved across every iteration of the loop body ($\{I\wedge B\}\,S\,\{I\}$), and upon termination ensures the postcondition ($I\wedge\lnot B\Rightarrow Q$).
Invariant synthesis amounts to generating a logical summary $I$ that is both correct, meaning it can be verified, and strong, meaning it enables verification of the final assertion. Weak but correct invariants contribute little, leaving most of the reasoning to the verifier, whereas strong invariants narrow the search space of program states, reduce solver effort, and yield substantial speedups.
3.2 Verifier-Based Decision Procedure
We formalize our verifier-based procedure for assessing candidate invariants. Let $P$ denote a program. A property is written as $p=\langle\varphi,\ell\rangle$, where $\varphi$ is a state predicate and $\ell$ is a program location. For a finite set $A$ of properties, let $\mathrm{Asm}(P,A)$ be the program obtained from $P$ by inserting assume statements for all elements of $A$. An execution of $\mathrm{Asm}(P,A)$ that reaches a location where an assumption is violated terminates immediately. We write $P\models_{A}p$ to indicate that all executions of $\mathrm{Asm}(P,A)$ satisfy the assertion $p$. The notation $P\models p$ abbreviates $P\models_{\varnothing}p$. Since assumptions restrict behaviors, if $P\not\models_{A}p$ for some $A$, then necessarily $P\not\models p$.
We assume access to a verifier
$V(P,A,p)\in\{\mathbf{T},\mathbf{F},\mathbf{U}\},$
which returns either $\mathbf{T}$ (proved), $\mathbf{F}$ (refuted), or $\mathbf{U}$ (inconclusive). The verifier is required to be sound on conclusive outcomes:
$V(P,A,p)=\mathbf{T}\;\Rightarrow\;P\models_{A}p,\qquad V(P,A,p)=\mathbf{F}\;\Rightarrow\;P\not\models_{A}p.$
No completeness is assumed for $\mathbf{U}$, which may arise from timeouts or incompleteness of the underlying verifier.
The verification task specifies a target property $p^{\star}=\langle\varphi^{\star},\ell^{\star}\rangle$. Given $P$ and $p^{\star}$, a large language model proposes a candidate invariant $q=\langle\psi,\ell\rangle$, typically at a loop header. To evaluate the utility of $q$, the procedure issues two verifier queries:
$d_{a}:=V(P,\varnothing,q)\quad\text{(checking whether $q$ is a correct predicate),}$
[…]
An example of the two verifier queries are given in Figure˜1. The outcome of the procedure is expressed as a judgment
$P\;\Rightarrow\;\langle p^{\star},q\rangle\;\Downarrow\;d\qquad\text{with}\quad d\in\{\mathbf{T},\mathbf{F},\mathbf{U}\}.$
The interpretation is as follows: if the judgment yields $\mathbf{T}$, then $p^{\star}$ is established on $P$; if it yields $\mathbf{F}$, then $p^{\star}$ is refuted; and if it yields $\mathbf{U}$, the attempt is inconclusive.
The inference rules defining this judgment are given below. Each rule specifies one possible derivation of the outcome, depending only on the responses of the verifier.
[…]
Rule DEC-FALSE captures short-circuit refutation: if the goal fails even in the restricted program $\mathrm{Asm}(P,\{q\})$, then it is false on the original program $P$. Rule DEC-PROP implements the prove-then-use strategy: once the candidate invariant $q$ is established, the outcome is exactly the verifier’s answer on the goal under the assumption $q$, restricted to $d\in\{\mathbf{T},\mathbf{U}\}$ so as not to overlap with DEC-FALSE. Rule DEC-$\mathbf{U}$ gives explicit conditions for inconclusiveness: the goal is not refuted under $q$ and $q$ is not established as an invariant.
Theorem (Decision Soundness).
If $P\Rightarrow\langle p^{\star},q\rangle\Downarrow\mathbf{T}$ is derivable, then $P\models p^{\star}$. If $P\Rightarrow\langle p^{\star},q\rangle\Downarrow\mathbf{F}$ is derivable, then $P\not\models p^{\star}$.
The proof is provided in Section˜A.1. This theorem establishes that whenever the calculus derives a conclusive outcome, that outcome is correct. The inconclusive case $\mathbf{U}$ is deliberately conservative: it makes no claim about the truth or falsity of the property and safely captures verifier incompleteness or timeouts.
3.3 Implementation
We describe the implementation of our verifier-based evaluation framework. Given a program $P$ and a target property $p^{\star}$, the system must generate candidate invariants $q$ and evaluate them according to the decision procedure. When proposing an invariant $q=\langle\psi,\ell\rangle$, the model selects a program location $\ell$ and supplies the corresponding predicate $\psi$.
Syntactic Validation.
Before invoking the verifier, we apply syntactic checks to the generated predicate $\psi$. These checks ensure that $\psi$ can be safely interpreted as a state predicate and that its insertion as an assumption does not alter the program state. For instance, expressions that update variables (e.g., a += 1) are rejected. Only Boolean conditions over the program state are accepted.
Parallel Evaluation. For each candidate $q$, the procedure issues two verifier queries, namely $d_{a}=V(P,\varnothing,q)$ to check whether $q$ is an invariant and $d_{b}=V(P,\{q\},p^{\star})$ to check whether the target holds under the assumption $q$. These queries are executed in parallel in our implementation, which reduces latency and enables speedup when the proposed invariant is useful for verification. The final outcome is then derived exactly according to the decision calculus.
4 Experimental Setup
Dataset Split Avg. #Lines #Instances
Evaluation Easy 51 113
Hard 62 113
Training 42 3589
Table 1: Dataset statistics of InvBench.
Dataset from SV-COMP.
We construct our benchmark from SV-COMP (Beyer & Strejček, 2025), a standard competition in software verification, focusing on problems that require loop invariant synthesis. We collect a pool of 899 instances and run the state-of-the-art non-LLM verifier UAutomizer (Schüssele et al., 2024) with a 600-second timeout to record the solving time for each. Based on this runtime, we classify instances into an easy split (solved within 30 seconds) and a hard split (solved between 30 and 600 seconds). From this pool, we randomly sample 113 problems from each split, resulting in 226 instances in the final evaluation set. The selection process is fully automated with no manual cherry-picking.
Synthetic Dataset.
For LLM fine-tuning, we construct a synthetic corpus of verification problems using GPT-4o. Seed programs supplied in the prompt are selected to ensure no overlap with the 226 instances in the evaluation set, to avoid data leakage between training and evaluation. Each synthetic program is analyzed by UAutomizer, and we retain only those whose assertions are successfully proved. The invariants extracted from UAutomizer ’s execution logs constitute our final fine-tuning dataset, which contains 3589 problems paired with their invariants.
Metrics.
All speedup-related measurements are reported relative to the state-of-the-art non-LLM-based solver UAutomizer (Schüssele et al., 2024), which serves as the baseline solver. We evaluate LLMs along two dimensions: the correctness of the generated invariants and the performance improvements they provide. Correctness is judged by the decision procedure formalized in Section˜3 with a timeout set to the problem’s original solving time by UAutomizer. For comparisons between UAutomizer and other tools, we report the number of solved instances under varying time budgets. We include the model’s token generation time in all evaluations.
Models.
We benchmark Claude models from Anthropic, GPT models from OpenAI, and the Qwen family of models (Hui et al., 2024; Yang et al., 2025).
Hardware and OS.
Experiments were conducted on a server with Intel Xeon Platinum 8275CL CPUs (96 cores), 8 NVIDIA A100 GPUs, and 1.1 TB of memory, running Ubuntu 22.04.
5 Results
5.1 Results of LLMs
[…]
Qwen2.5-72B 4.4% 0.9% 1.20× 1.00×
gpt-oss-120b 8.8% 5.3% 1.10× 1.01×
claude-sonnet-4 15.0% 8.8% 1.06× 1.01×
Qwen3-Coder-480B 14.2% 8.0% 1.09× 1.01×
claude-opus-4.1 18.6% 8.8% 1.23× 1.02×
gpt-5 37.2% 26.5% 1.24× 1.06×
o3 39.8% 28.3% 1.37× 1.09×
[…]
gpt-5 11.5% 0% 1.00× 1.00×
Qwen2.5-72B 12.4% 0% 1.00× 1.00×
o3 29.2% 0% 1.00× 1.00×
Qwen3-Coder-480B 15.9% 0% 1.00× 1.00×
claude-sonnet-4 13.3% 0.9% 3.35× 1.01×
claude-opus-4.1 14.2% 0.9% 2.96× 1.02×
gpt-oss-120b 11.5% 0.9% 29.57× 1.03×
Table 3: InvBench-Hard: Results on the InvBench dataset (hard split) across models.
We report the performance of different LLMs on the easy and hard splits of InvBench in Table˜2 and Table˜3, respectively. On the easy split shown in Table˜2, o3 achieves the strongest results, with a 1.37× speedup over UAutomizer on 28.3% of problems and an overall 1.09× average speedup. These results indicate that while state-of-the-art LLMs generally struggle to synthesize correct or sufficiently strong invariants to accelerate program verification, the strongest model o3 can still yield non-trivial improvements on a meaningful fraction of problems.
Table˜3 reports results on the hard split of InvBench. In this setting, LLMs show only negligible improvement over UAutomizer. The only noteworthy exception is gpt-oss-120b, which produces an invariant for one problem that delivers a 29.57× speedup. However, overall, such speedup cases are exceedingly rare.
There are three main takeaways from the results. First, generating strong invariants that yield performance speedups is substantially more difficult than merely producing correct invariants, as reflected in the large gap between the percentage of correct invariants and the percentage of speedups. Second, model capability is a key factor, as demonstrated by the sharp differences in speedups reported in Table˜2. Third, LLMs remain far from fully addressing the task of invariant synthesis, leaving considerable room for future progress, as shown in Table˜3.
5.2 Results of Verifiers on InvBench
We compare the state-of-the-art non-LLM-based tool UAutomizer (Schüssele et al., 2024) with other LLM-based verifiers on both the easy split and the hard split of InvBench.
On the easy split of InvBench, where UAutomizer solves all 113 instances within 30 seconds, it clearly surpasses all LLM-based verifiers. LaM4Inv (Wu et al., 2024a) and Loopy (Kamath et al., 2023) fail to solve any instance within this time limit, while LEMUR (Wu et al., 2024b) solves only 55 instances. These results indicate that none of the LLM-based verifiers provide any advantage over UAutomizer on the easy split.
Table˜4 presents the comparison on the hard split of InvBench. UAutomizer still delivers the best performance across almost all timeouts. UAutomizer consistently and significantly outperforms LaM4Inv and Loopy. In particular, LaM4Inv and Loopy solve fewer than half as many instances as UAutomizer across all timeouts. Notably, within 10 seconds, LEMUR solves more problems than UAutomizer, suggesting that it can accelerate some of the instances. Nevertheless, UAutomizer still outperforms LEMUR with longer timeouts, indicating that although LEMUR can offer early solving advantages on certain challenging instances, its overall capability is still limited compared to UAutomizer.
Given the consistently poor performance of LaM4Inv, which often either fails or times out, we conducted a manual investigation. Our analysis shows that LaM4Inv fails on some examples when applied to external datasets beyond those used in its customized evaluation. However, the preprocessing steps or manual annotations required by the tool are not documented. The official repository does not provide preprocessing scripts, and our attempt to contact the authors has not received a response. For a direct comparison between UAutomizer and LaM4Inv, we also refer readers to Section˜5.3, where we evaluate UAutomizer on the dataset released by LaM4Inv.
5.3 Results of UAutomizer on Prior Datasets
As shown in Table 5, UAutomizer, the state-of-the-art non-LLM-based verifier, consistently solves more instances than LaM4Inv (Wu et al., 2024a) and Loopy (Kamath et al., 2023), two representative LLM-based tools, on their respective customized datasets. This highlights that prior work omitted an important baseline comparison against UAutomizer.
LEMUR (Wu et al., 2024b) is the only tool that surpasses UAutomizer. However, this is largely explained by its dataset construction: LEMUR evaluates only on problems that UAutomizer cannot solve within 600 seconds. To assess whether this advantage generalizes beyond its own curated benchmark, we conduct an additional analysis presented in Section˜6.
While LaM4Inv and Loopy consistently underperform UAutomizer in terms of the total number of solved instances across time budgets, they nevertheless offer complementary benefits, which we discuss in Section˜A.2.
We further investigate the distributional differences between datasets. At the 600-second timeout, the performance of prior LLM-based verifiers approaches that of UAutomizer, with LaM4Inv even matching it exactly. In contrast, on InvBench (see Section˜5.2), all LLM-based verifiers perform poorly at 600 seconds, suggesting a substantial shift in distribution compared to the customized datasets used in prior work. Our analysis confirms this: programs in InvBench are significantly longer, averaging 62 lines of code in the hard split, compared to only 22 for LaM4Inv, 23 for LEMUR, and 27 for Loopy. Manual inspection shows that InvBench contains features such as multiple loops, functions, arrays, and pointers, which are largely absent from prior datasets. This makes InvBench a more challenging and realistic benchmark that better distinguishes solver performance.
We also note that different LLM-based verification frameworks are built on different base solvers: LEMUR is built on UAutomizer, Loopy on Frama-C (Cuoq et al., 2012), and LaM4Inv on ESBMC (Gadelha et al., 2018). Given the strength of UAutomizer, we believe future work should place greater emphasis on developing LLM-based verifiers atop state-of-the-art solvers such as UAutomizer and ensure that comparisons against it are not omitted.
Verifier Solved Instances under Different Timeouts
10s 100s 300s 600s
LaM4Inv (Wu et al., 2024a) 0 0 0 0
Loopy (Kamath et al., 2023) 0 2 18 50
LEMUR (Wu et al., 2024b) 32 71 73 73
UAutomizer 0 73 95 113
Table 4: Comparison of solved instances on InvBench-Hard by different verifiers under varying timeout budgets.
Verifier Total Instances Solved Instances under Different Timeouts
10s 100s 300s 600s
LaM4Inv (Wu et al., 2024a) 316 144 286 295 299
UAutomizer 316 299 299 299 299
Loopy (Kamath et al., 2023) 469 0 133 353 403
UAutomizer 469 372 403 411 413
LEMUR (Wu et al., 2024b) 47 2 8 16 19
UAutomizer 47 0 0 0 0
Table 5: Comparison of prior LLM-based verifiers and UAutomizer on their own customized dataset under different timeout budgets.
5.4 Performance Gains from Invariants Generated by UAutomizer
To quantify the potential speedup achievable when providing correct and strong invariants to solvers, we conduct a study using UAutomizer. We extract the invariants identified by UAutomizer and then measure the resulting speedup when they are supplied to the verifier. On a random sample of 100 problems from our training set, we observe an overall average speedup of 1.86×. This indicates that the invariants discovered by UAutomizer are sufficiently strong to accelerate verification, and we leverage them for fine-tuning LLMs in Section˜5.5.
5.5 Effectiveness of Fine-Tuning
For each synthetically generated problem in the training set, we run UAutomizer to produce invariants, yielding a training dataset of 3589 problems with associated invariants. We then perform supervised fine-tuning using LoRA for 3 epochs.
[…]
Qwen2.5-72B (base) 4.4% 0.9% 1.20× 1.00×
Qwen2.5-72B (fine-tuned) 32.7% 26.5% 1.13× 1.03×
Qwen3-Coder-480B (base) 14.2% 8.0% 1.09× 1.01×
Qwen3-Coder-480B (fine-tuned) 40.7% 29.2% 1.29× 1.08×
Table 6: Improvement of supervised fine-tuning on InvBench-Easy.
Table˜6 reports results on the easy split of InvBench. Fine-tuning leads to substantial gains in both invariant correctness and runtime speedup. Qwen3-Coder-480B shows the most pronounced improvements: the proportion of correct invariants increases from 14.2% in the base model to 40.7%, and the percentage of instances with speedups rises from 8% to 29.2%. For Qwen2.5-72B, the conditional speedup decreases after fine-tuning, but this is because the baseline conditional speedup was driven by a single case, whereas fine-tuning yields a larger number of cases with speedups. We also note that on the hard split, neither of the fine-tuned models shows improvement. These results suggest that more training efforts are required to achieve gains on harder instances.
5.6 Effectiveness of Best-of-N Sampling
[…]
claude-opus-4.1 1 18.6% 8.8% 1.23× 1.02×
claude-opus-4.1 16 37.2% 23.0% 1.15× 1.03×
claude-sonnet-4 1 15.0% 8.8% 1.06× 1.01×
claude-sonnet-4 16 36.3% 22.1% 1.20× 1.04×
o3 1 39.8% 28.3% 1.37× 1.09×
o3 16 50.4% 35.4% 1.39× 1.12×
Table 7: Results of Best-of-N sampling (N = 16) on InvBench-Easy.
We also evaluated whether repeated sampling improves LLM performance in invariant synthesis. For each problem, we generated 16 samples at a temperature of 0.7, removed duplicate invariants, and verified all candidates in parallel. We then select the sample with the shortest verification time and kill all other samples. Table 7 summarizes the results on InvBench’s easy split, demonstrating that repeated sampling leads to consistent gains. For example, o3 improves from 39.8% to 50.4% correct invariants, while claude-opus-4.1 nearly doubles its accuracy from 18.6% to 37.2%.
On the hard split, for claude-opus-4.1, with 16 samples, the percentage of instances with correct invariants increases from 14.2% to 15.9%, the percentage of instances with speedup from 0.9% to 2.7%, and the overall average speedup from 1.02× to 1.03×. Similarly, for claude-sonnet-4, the percentage of instances with correct invariants increases from 13.3% to 20.4%, the percentage of instances with speedup from 0.9% to 2.7%, conditional average speedup from 3.35× to 6.12×, and the overall average speedup from 1.01× to 1.07×. These improvements suggest that best-of-N sampling is a promising technique for LLMs to improve invariant synthesis performance.
6 Discussion
Generalizability of LEMUR.
LEMUR (Wu et al., 2024b) reports the best results on its customized dataset, as it targets instances that UAutomizer cannot solve within 600 seconds. To test whether this advantage generalizes, we sampled 50 unsolved instances and found that LEMUR solved 12 within the same timeout of 600 seconds. This suggests that LEMUR’s gains are not solely due to benchmark design but reflect a genuine advantage on harder problems.
Inference Overhead of Models.
Our evaluation includes LLM serving time as a realistic measure of end-to-end performance. Since the goal is to accelerate verification, inference overhead must be considered alongside solver runtime. Future research may explore how to balance the quality of generated invariants with the inference cost of producing them.
7 Conclusion
This work introduced InvBench, a principled framework for evaluating the capabilities of LLMs in invariant synthesis. Our approach employs a verifier-based decision procedure with a formal soundness guarantee and assesses not only correctness but also the speedups that invariants contribute to program verification. Using a benchmark of 226 instances, we conducted a comparison across state-of-the-art LLMs, existing LLM-based verifiers, and the traditional solver UAutomizer. The results show that although LLM-based verifiers represent a promising direction, they do not yet offer significant advantages over non-LLM-based approaches. Model capability proves to be a critical factor, and our benchmark remains an open challenge for current LLMs. At the same time, we demonstrated that supervised fine-tuning and Best-of-N sampling can improve model performance in accelerating verification.
LLM Usage
LLMs are the subject of this study. We additionally used them for polishing the writing.
Appendix A Appendix
A.1 Proof
Theorem (Decision Soundness).
If $P\Rightarrow\langle p^{\star},q\rangle\Downarrow\mathbf{T}$ is derivable, then $P\models p^{\star}$. If $P\Rightarrow\langle p^{\star},q\rangle\Downarrow\mathbf{F}$ is derivable, then $P\not\models p^{\star}$.
Proof.
For the $\mathbf{T}$ case, the final rule must be DEC-PROP with $d=\mathbf{T}$. The premises yield $V(P,\varnothing,q)=\mathbf{T}$ and $V(P,\{q\},p^{\star})=\mathbf{T}$. By verifier soundness, $P\models q$ and $P\models_{\{q\}}p^{\star}$. Since $q$ holds on all executions of $P$, introducing the assumption $\{q\}$ does not remove executions relevant to $p^{\star}$; thus $P\models p^{\star}$.
For the $\mathbf{F}$ case, the final rule must be DEC-FALSE. Its premise $V(P,\{q\},p^{\star})=\mathbf{F}$ implies $P\not\models_{\{q\}}p^{\star}$ by soundness. Assumptions restrict behaviors; hence, a violation under assumptions entails a violation without them, yielding $P\not\models p^{\star}$. ∎
A.2 Complementary Results of LLM-based Solvers
As shown in Table˜A1, all three LLM-based tools solve instances that remain unsolved by UAutomizer within the 600-second budget. LaM4Inv and Loopy each contribute additional solved cases, and LEMUR is able to handle 19 problems that UAutomizer cannot solve at all, underscoring the complementary strengths of LLM-based approaches.
Verifier Solved Instances $\Delta$ Solved
LaM4Inv 299 13
Loopy 403 40
LEMUR 19 19
Table A1: LLM-based verifiers complement UAutomizer by solving problems beyond its reach. We report the total numbers of solved problems, and $\Delta$ Solved” is the number of instances uniquely solved that are unsolved by UAutomizer.

### v2

@@
-InvBench: Can LLMs Accelerate Program Verification with Invariant Synthesis?
+Quokka: Accelerating Program Verification with LLMs via Invariant Synthesis
 Abstract
 Program verification relies on loop invariants, yet automatically discovering strong invariants remains a long-standing challenge.
-We introduce a principled framework for evaluating LLMs on invariant synthesis.
-Our approach uses a verifier-based decision procedure with a formal soundness guarantee and assesses not only correctness but also the speedup that invariants provide in verification.
-We evaluate 7 state-of-the-art LLMs, and existing LLM-based verifiers against the traditional solver UAutomizer.
-While LLM-based verifiers represent a promising direction, they do not yet offer a significant advantage over UAutomizer.
-Model capability also proves critical, as shown by sharp differences in speedups across models, and our benchmark remains an open challenge for current LLMs.
-Finally, we show that supervised fine-tuning and Best-of-N sampling can improve performance: fine-tuning on 3589 instances raises the percentage of speedup cases for Qwen3-Coder-480B from 8% to 29.2%, and Best-of-N sampling with N=16 improves Claude-sonnet-4 from 8.8% to 22.1%.
+We investigate whether large language models (LLMs) can accelerate program verification by generating useful loop invariants.
+We introduce Quokka, a first-order and effective framework for LLM-based invariant synthesis that provides sound evaluation while achieving state-of-the-art speedup results.
+Unlike prior work that designs complex, highly customized algorithms, Quokka employs a simple and principled verification procedure.
+We construct a benchmark of 866 instances and evaluate 9 state-of-the-art LLMs across multiple model families.
+Our results show that Quokka consistently outperforms all prior LLM-based verifiers: achieving speedups of at least 1.2× on 81 instances compared to 39 instances for the previous best approach.
+We further demonstrate that supervised fine-tuning and Best-of-N sampling can yield measurable improvements in accelerating verification.
+Machine Learning, ICML
 1 Introduction
@@
 Traditional approaches include constraint solving (Colon et al., 2003; Gupta et al., 2009), dynamic analysis (Le et al., 2019), etc.
-Since invariant discovery is undecidable in general, researchers have tried a variety of learning-based methods (Li et al., 2017; Ezudheen et al., 2018).
-Building on this progression, the strong capabilities of LLMs in code generation and program reasoning (Austin et al., 2021; Chen et al., 2021; Wei et al., 2025b; c) naturally motivate a systematic evaluation of their potential for invariant discovery.
-Pei et al.
-(2023) is the first work to evaluate the capabilities of LLMs in invariant generation.
-However, their methodology considers only correctness and does not assess how strong the generated invariants are.
-As a result, LLMs may generate correct invariants that perform well under their evaluation metric but provide no benefit for accelerating the verification process in real-world settings.
-Furthermore, their notion of correctness is not based on formal verification.
-Instead, it is determined by direct comparison with invariants generated by an existing tool, namely Daikon (Ernst et al., 2007).
-Daikon is a dynamic analysis tool whose invariants are not guaranteed to be sound, since they are inferred from observed test executions rather than proven across all possible executions.
-As a result, the ground-truth invariants themselves may be incorrect.
-Moreover, directly comparing LLM-generated invariants with Daikon’s output can lead to rejecting many correct invariants.
-For example, if an LLM proposes $a>0$ for an integer $a$ while Daikon reports $a\geq 1$, the evaluation in Pei et al.
-(2023) would incorrectly classify the LLM’s result as wrong, even though the two are equivalent.
-Hence, prior work’s evaluation methodology cannot reliably capture the correctness of LLM-generated invariants, let alone how useful they are for verification.
-A series of follow-up works inspired by Pei et al.
-(2023) have proposed LLM-based verifiers (Wu et al., 2024b; a; Kamath et al., 2024).
+Since invariant discovery is challenging, researchers have tried a variety of learning-based methods (Li et al., 2017; Ezudheen et al., 2018).
+Building on this progression, the strong capabilities of LLMs in code generation and program reasoning (Austin et al., 2021; Chen et al., 2021; Wei et al., 2025b) naturally motivate a systematic evaluation of their potential for invariant discovery.
+The first work (Pei et al., 2023) to evaluate LLMs for invariant generation has two key limitations.
+One limitation is that it only checks correctness without assessing whether invariants are strong enough to accelerate verification.
+Another limitation is that correctness is determined by directly comparing against Daikon (Ernst et al., 2007), a dynamic analysis tool that infers invariants from test executions rather than formal verification.
+This approach is unsound: Daikon’s invariants may themselves be incorrect.
+Moreover, semantically equivalent invariants (e.g., $a>0$ vs.
+$a\geq 1$ for integers) can be incorrectly rejected as wrong.
+Consequently, prior work cannot reliably evaluate either the correctness or the practical utility of LLM-generated invariants.
+A series of follow-up works have proposed LLM-based verifiers.
 Instead of evaluating LLMs in isolation, these efforts develop verification frameworks powered by LLMs.
-However, the evaluation of such tools suffers from two key limitations: (1) each work introduces a customized dataset and reports results only on that dataset, making cross-comparison impossible; and (2) many omit direct comparison with state-of-the-art traditional verifiers, underscoring the need for evaluation against established baselines.
-Figure 1: Illustration of InvBench’s evaluation pipeline.
+This line of work predominantly adopts complex, sequential, and iterative processes with highly customized algorithms to handle LLM-generated invariants.
+For instance, LaM4Inv (Wu et al., 2024a) uses a “query-filter-reassemble” strategy that filters and combines model-generated predicates through complex logical operations (conjunctions or disjunctions) to construct valid loop invariants, rather than using them directly.
+LOOPY (Kamath et al., 2024) employs a customized Houdini algorithm with iterative refinement to filter candidate invariants.
+LEMUR (Wu et al., 2024b) develops a backtracking and sequential algorithm.
+While these approaches demonstrate the potential of LLM-based verification, their complexity raises a question: can we design a simpler, principled, and more effective decision procedure for LLM-based verifiers?
+Figure 1: Illustration of Quokka’s evaluation pipeline.
 The LLM proposes an invariant by specifying a program location and predicate (e.g., location B with x % 7 == 3).
-The verification procedure then incorporates this invariant to prove the property x != 145 using two verifier queries, and we measure the resulting speedup relative to a baseline without LLM assistance.
-Our work, InvBench, introduces a principled methodology for evaluating the capabilities of LLMs in invariant synthesis.
-Instead of checking whether LLMs reproduce invariants discovered by other tools, we employ a verifier-based decision procedure that directly determines the correctness of LLM-generated invariants, and we prove this procedure to be sound.
-We formalize the methodology as a proof calculus, providing a rigorous foundation for invariant evaluation.
-Unlike previous work that focuses on optimizing verification performance by designing tailored interactive protocols between solvers and LLMs (Chakraborty et al., 2023; Wu et al., 2024b; Kamath et al., 2024; Wu et al., 2024a), our goal is to develop a simple, first-order verification procedure suitable for evaluating the invariant generation capability of LLMs.
-Our formalization provides such a framework, offering a general solution for invariant evaluation.
-Since the purpose of invariant synthesis is to accelerate verification, invariants that are too weak to aid verification or too difficult to verify as correct offer little practical value.
-To capture this, we evaluate the invariants by measuring the speedup they provide in the overall verification process.
-To support comparison across solvers and LLMs, we construct a dataset of 226 instances derived from the most recent edition of the software verification competition SV-Comp (Beyer & Strejček, 2025) and use it to evaluate multiple LLM-based verifiers on this common benchmark.
-In addition, we assess the state-of-the-art traditional (i.e., non-LLM-based) verifier UAutomizer (Schüssele et al., 2024) both on our dataset and on each of the customized datasets introduced in prior work (Wu et al., 2024b; Kamath et al., 2024; Wu et al., 2024a).
-UAutomizer consistently outperforms prior LLM-based verifiers on both our dataset and their customized datasets across all settings, with the only exception of LEMUR (Wu et al., 2024b), which was specifically designed for problems that UAutomizer fails to solve.
-These results suggest that while LLM-based verifiers are a promising direction, they do not yet offer a significant advantage over non-LLM-based approaches in general.
-Our findings further highlight the importance of using established datasets and baselines for meaningful progress in the community.
-They also highlight the role of the underlying solver, suggesting that base solver choice can influence outcomes.
-In addition, we evaluate 7 state-of-the-art LLMs.
-The results indicate that model capability is critical, as shown by the sharp differences in achieved speedups.
-Nonetheless, LLMs remain largely unable to solve the hard subset of InvBench, underscoring that our dataset presents an open challenge for current LLMs.
-We also investigate the extent to which supervised fine-tuning and Best-of-N sampling can enhance model performance in accelerating verification.
-To this end, we construct a fine-tuning dataset of 3589 instances and show that training on this dataset raises the percentage of speedup cases for Qwen3-Coder-480B from 8% to 29.2%.
-Similarly, Best-of-N sampling with N = 16 improves Claude-sonnet-4 from 8.8% to 22.1%.
+The verification procedure then incorporates this invariant to prove the property x != 700 using two verifier queries, and we measure the resulting speedup relative to a baseline without LLM assistance.
+In this work, we introduce Quokka, a framework for LLM-based invariant synthesis that accelerates program verification.
+Quokka provides a sound approach to evaluate LLMs’ capabilities in generating invariants while achieving state-of-the-art speedup results compared with prior LLM-based verifiers.
+We formalize our methodology as a proof calculus, providing a rigorous foundation for building LLM-based verifiers.
+Instead of checking the correctness of invariants based on string-matching, we adopt a verifier-based approach that directly queries the underlying solver.
+Unlike previous work that designs complex, highly customized algorithms, Quokka employs a simple and principled verification procedure.
+Our approach enables both soundness guarantees and precise measurement of performance speedup, allowing us to systematically evaluate whether LLM-generated invariants can accelerate verification in practice.
+To support systematic comparison across solvers and LLMs, we construct a dataset of 866 instances derived from the most recent edition of the software verification competition SV-COMP (Beyer and Strejček, 2025).
+To the best of our knowledge, this is the largest evaluation dataset for LLM-based verifiers to date, compared with datasets used in previous work.
+Quokka consistently outperforms all prior LLM-based verifiers.
+When comparing against LEMUR (the previous best LLM-based verifier with the same underlying solver), Quokka achieves speedups of at least 1.2× (i.e., verification completes at least 20% faster than the baseline solver) on 81 instances compared to LEMUR’s 39 instances.
+For more substantial speedups of at least 2.0× (i.e., verification completes at least twice as fast), Quokka reaches 51 instances compared to LEMUR’s 22 instances.
+We evaluate 9 state-of-the-art LLMs, including GPT models from OpenAI, Claude models from Anthropic, and models from the Qwen and LLaMA families.
+Our experiments reveal that gpt-5.2 achieves the strongest speedup results among all evaluated models.
+Beyond model evaluation, we investigate techniques to enhance invariant generation capabilities.
+We construct a training dataset of 3589 instances using a verifier-based filtering approach that ensures data quality.
+Our experiments demonstrate that both supervised fine-tuning and Best-of-N sampling yield measurable improvements in accelerating verification.
+Specifically, Best-of-8 sampling achieves speedups of at least 1.2× on 99 instances, representing a 22% improvement over single-sample generation, which achieves such speedups on only 81 instances.
 In summary, our contributions are as follows:
-  • We propose a verification procedure for evaluating the invariant generation capabilities of LLMs, assessing both correctness and their effectiveness in accelerating verification.
-  • We evaluate 7 state-of-the-art LLMs, and provide comparisons between existing LLM-based verifiers and the state-of-the-art non-LLM-based solver UAutomizer.
-  • We construct a dataset of 3589 instances for training. We demonstrate that both supervised fine-tuning and Best-of-N sampling can improve model performance in accelerating verification.
+  • We introduce a decision procedure for LLM-based invariant synthesis that provides sound evaluation and achieves the best speedup results compared with prior LLM-based verifiers.
+  • We construct a benchmark of 866 instances, and conduct an evaluation of 9 state-of-the-art LLMs across multiple model families.
+  • We construct training data and demonstrate that both supervised fine-tuning and Best-of-N sampling yield measurable improvements in accelerating verification.
 2 Related Work
 Traditional Methods for Program Invariant Generation.
-A long line of research has explored invariant synthesis using traditional techniques without machine learning, including model checking (Flanagan & Qadeer, 2002; Lahiri & Bryant, 2007; Hojjat & Rümmer, 2018; Vediramana Krishnan et al., 2024), abstract interpretation (Karr, 1976; Cousot & Cousot, 1977; Cousot & Halbwachs, 1978; Cousot & Cousot, 1979), constraint solving (Gulwani et al., 2009; Gupta et al., 2009), Craig interpolation (Jhala & McMillan, 2006; McMillan, 2010), and syntax-guided synthesis (Fedyukovich & Bodík, 2018).
+A long line of research has explored invariant synthesis using traditional techniques without machine learning, including model checking (Flanagan and Qadeer, 2002; Lahiri and Bryant, 2007; Hojjat and Rümmer, 2018; Vediramana Krishnan et al., 2024), abstract interpretation (Karr, 1976; Cousot and Cousot, 1977; Cousot and Halbwachs, 1978; Cousot and Cousot, 1979), constraint solving (Gulwani et al., 2009; Gupta et al., 2009), Craig interpolation (Jhala and McMillan, 2006; McMillan, 2010), and syntax-guided synthesis (Fedyukovich and Bodík, 2018).
 Prior work evaluating LLM-generated invariants (Pei et al., 2023) has relied on Daikon (Ernst et al., 2007), a tool for dynamic invariant detection (Echenim et al., 2019; Le et al., 2019).
@@
 Our approach instead employs a verifier-based decision procedure relying on UAutomizer (Schüssele et al., 2024) that ensures soundness.
-Learning-Based Method for Invariant Generation
-Machine learning based techniques have been widely adopted in invariant synthesis, including decision tree (Garg et al., 2014; 2016; Ezudheen et al., 2018; Riley & Fedyukovich, 2022; Xu et al., 2020), support vector machine (Li et al., 2017; Sharma et al., 2012), reinforcement learning (Si et al., 2018; Yu et al., 2023), and others (Sharma et al., 2013; Ryan et al., 2019; Yao et al., 2020).
-More recently, large language models have demonstrated strong capabilities in reasoning about code and logic (Wei et al., 2025a; b; d), giving rise to a series of work that explore using LLMs for finding invariants.
-Pei et al.
-(2023) is the first pioneering work that evaluates LLMs’ capabilities in finding invariants, but it is not a sound evaluation.
-Chakraborty et al.
-(2023) proposes ranking the invariants generated by LLMs.
-LaM4Inv (Wu et al., 2024a) proposes a “query-filter-reassemble” strategy to leverage the generation power of LLMs, LEMUR (Wu et al., 2024b) proposes a sound algorithm with back-tracking.
-Loopy (Kamath et al., 2023) builds an LLM-based pipeline that combines clause sampling, Houdini filtering, and error-guided repair.
-Liu et al.
-(2024) proposes a new dataset including problems from courses and real-world systems.
-Our work explores fine-tuning and the best-of-N sampling method to improve LLMs’ capabilities.
+Learning-Based Method for Invariant Generation.
+Machine learning based techniques have been widely adopted in invariant synthesis, including decision tree (Garg et al., 2014, 2016; Ezudheen et al., 2018; Riley and Fedyukovich, 2022; Xu et al., 2020), support vector machine (Li et al., 2017; Sharma et al., 2012), reinforcement learning (Si et al., 2018; Yu et al., 2023), and others (Sharma et al., 2013; Ryan et al., 2019; Yao et al., 2020).
+More recently, large language models have demonstrated strong capabilities in reasoning about code and logic (Wei et al., 2025a, b, c), giving rise to a series of work that explore using LLMs for finding invariants.
+(Pei et al., 2023) is the first pioneering work that evaluates LLMs’ capabilities in finding invariants, but it is not a sound evaluation.
+Various techniques have been proposed to couple LLMs with symbolic solvers, including ranking LLM-generated invariants (Chakraborty et al., 2023), the “query-filter-reassemble” strategy of LaM4Inv (Wu et al., 2024a), the back-tracking algorithm in LEMUR (Wu et al., 2024b), and Loopy’s integration of the classic Houdini algorithm (Kamath et al., 2024).
+On the dataset side, (Liu et al., 2024) introduces a rule-based method for constructing a fine-tuning corpus, which differs from our verifier-based approach.
+In contrast, our work provides a simple and sound evaluation procedure for assessing LLM-generated invariants and investigates how both fine-tuning and Best-of-N sampling can enhance LLM performance in invariant synthesis.
 3 Method
 3.1 Preliminary
-We formalize the task of loop invariant synthesis using standard Hoare logic Hoare (1969).
+We formalize the task of loop invariant synthesis using standard Hoare logic (Hoare, 1969).
 A Hoare triple $\{P\}\,S\,\{Q\}$ specifies that if the precondition $P$ holds before executing a statement $S$, then the postcondition $Q$ will hold after its execution.
@@
 which returns either $\mathbf{T}$ (proved), $\mathbf{F}$ (refuted), or $\mathbf{U}$ (inconclusive). The verifier is required to be sound on conclusive outcomes:
-$V(P,A,p)=\mathbf{T}\;\Rightarrow\;P\models_{A}p,\qquad V(P,A,p)=\mathbf{F}\;\Rightarrow\;P\not\models_{A}p.$
+$V(P,A,p)=\mathbf{T}\;\Rightarrow\;P\models_{A}p,$
+$V(P,A,p)=\mathbf{F}\;\Rightarrow\;P\not\models_{A}p.$
 No completeness is assumed for $\mathbf{U}$, which may arise from timeouts or incompleteness of the underlying verifier.
@@
 To evaluate the utility of $q$, the procedure issues two verifier queries:
-$d_{a}:=V(P,\varnothing,q)\quad\text{(checking whether $q$ is a correct predicate),}$
-[…]
-An example of the two verifier queries are given in Figure˜1. The outcome of the procedure is expressed as a judgment
+$d_{a}:=V(P,\varnothing,q)$
+$d_{b}:=V(P,\{q\},p^{\star})$
+The query $d_{a}$ verifies whether $q$ is a correct invariant, while the query $d_{b}$ determines whether the target property holds under the assumption that $q$ is true.
+An illustration of these two verifier queries is provided in Figure˜1: $d_{a}$ corresponds to assert(x % 7 == 3) (upper subfigure), and $d_{b}$ corresponds to verifying the final assertion with the assume statement inserted (lower subfigure).
+The outcome of the procedure is expressed as a judgment
 $P\;\Rightarrow\;\langle p^{\star},q\rangle\;\Downarrow\;d\qquad\text{with}\quad d\in\{\mathbf{T},\mathbf{F},\mathbf{U}\}.$
@@
 Rule DEC-$\mathbf{U}$ gives explicit conditions for inconclusiveness: the goal is not refuted under $q$ and $q$ is not established as an invariant.
-Theorem (Decision Soundness).
-If $P\Rightarrow\langle p^{\star},q\rangle\Downarrow\mathbf{T}$ is derivable, then $P\models p^{\star}$.
-If $P\Rightarrow\langle p^{\star},q\rangle\Downarrow\mathbf{F}$ is derivable, then $P\not\models p^{\star}$.
-The proof is provided in Section˜A.1.
-This theorem establishes that whenever the calculus derives a conclusive outcome, that outcome is correct.
-The inconclusive case $\mathbf{U}$ is deliberately conservative: it makes no claim about the truth or falsity of the property and safely captures verifier incompleteness or timeouts.
-3.3 Implementation
-We describe the implementation of our verifier-based evaluation framework.
-Given a program $P$ and a target property $p^{\star}$, the system must generate candidate invariants $q$ and evaluate them according to the decision procedure.
-When proposing an invariant $q=\langle\psi,\ell\rangle$, the model selects a program location $\ell$ and supplies the corresponding predicate $\psi$.
-Syntactic Validation.
-Before invoking the verifier, we apply syntactic checks to the generated predicate $\psi$.
-These checks ensure that $\psi$ can be safely interpreted as a state predicate and that its insertion as an assumption does not alter the program state.
-For instance, expressions that update variables (e.g., a += 1) are rejected.
-Only Boolean conditions over the program state are accepted.
-Parallel Evaluation.
-For each candidate $q$, the procedure issues two verifier queries, namely $d_{a}=V(P,\varnothing,q)$ to check whether $q$ is an invariant and $d_{b}=V(P,\{q\},p^{\star})$ to check whether the target holds under the assumption $q$.
-These queries are executed in parallel in our implementation, which reduces latency and enables speedup when the proposed invariant is useful for verification.
-The final outcome is then derived exactly according to the decision calculus.
-4 Experimental Setup
-Dataset Split Avg. #Lines #Instances
-Evaluation Easy 51 113
-Hard 62 113
-Training 42 3589
-Table 1: Dataset statistics of InvBench.
-Dataset from SV-COMP.
-We construct our benchmark from SV-COMP (Beyer & Strejček, 2025), a standard competition in software verification, focusing on problems that require loop invariant synthesis.
-We collect a pool of 899 instances and run the state-of-the-art non-LLM verifier UAutomizer (Schüssele et al., 2024) with a 600-second timeout to record the solving time for each.
-Based on this runtime, we classify instances into an easy split (solved within 30 seconds) and a hard split (solved between 30 and 600 seconds).
-From this pool, we randomly sample 113 problems from each split, resulting in 226 instances in the final evaluation set.
-The selection process is fully automated with no manual cherry-picking.
-Synthetic Dataset.
-For LLM fine-tuning, we construct a synthetic corpus of verification problems using GPT-4o.
-Seed programs supplied in the prompt are selected to ensure no overlap with the 226 instances in the evaluation set, to avoid data leakage between training and evaluation.
-Each synthetic program is analyzed by UAutomizer, and we retain only those whose assertions are successfully proved.
-The invariants extracted from UAutomizer ’s execution logs constitute our final fine-tuning dataset, which contains 3589 problems paired with their invariants.
-Metrics.
-All speedup-related measurements are reported relative to the state-of-the-art non-LLM-based solver UAutomizer (Schüssele et al., 2024), which serves as the baseline solver.
-We evaluate LLMs along two dimensions: the correctness of the generated invariants and the performance improvements they provide.
-Correctness is judged by the decision procedure formalized in Section˜3 with a timeout set to the problem’s original solving time by UAutomizer.
-For comparisons between UAutomizer and other tools, we report the number of solved instances under varying time budgets.
-We include the model’s token generation time in all evaluations.
-Models.
-We benchmark Claude models from Anthropic, GPT models from OpenAI, and the Qwen family of models (Hui et al., 2024; Yang et al., 2025).
-Hardware and OS.
-Experiments were conducted on a server with Intel Xeon Platinum 8275CL CPUs (96 cores), 8 NVIDIA A100 GPUs, and 1.1 TB of memory, running Ubuntu 22.04.
-5 Results
-5.1 Results of LLMs
-[…]
-Qwen2.5-72B 4.4% 0.9% 1.20× 1.00×
-gpt-oss-120b 8.8% 5.3% 1.10× 1.01×
-claude-sonnet-4 15.0% 8.8% 1.06× 1.01×
-Qwen3-Coder-480B 14.2% 8.0% 1.09× 1.01×
-claude-opus-4.1 18.6% 8.8% 1.23× 1.02×
-gpt-5 37.2% 26.5% 1.24× 1.06×
-o3 39.8% 28.3% 1.37× 1.09×
-[…]
-[…]
-gpt-5 11.5% 0% 1.00× 1.00×
-Qwen2.5-72B 12.4% 0% 1.00× 1.00×
-o3 29.2% 0% 1.00× 1.00×
-Qwen3-Coder-480B 15.9% 0% 1.00× 1.00×
-claude-sonnet-4 13.3% 0.9% 3.35× 1.01×
-claude-opus-4.1 14.2% 0.9% 2.96× 1.02×
-gpt-oss-120b 11.5% 0.9% 29.57× 1.03×
-Table 3: InvBench-Hard: Results on the InvBench dataset (hard split) across models.
-We report the performance of different LLMs on the easy and hard splits of InvBench in Table˜2 and Table˜3, respectively.
-On the easy split shown in Table˜2, o3 achieves the strongest results, with a 1.37× speedup over UAutomizer on 28.3% of problems and an overall 1.09× average speedup.
-These results indicate that while state-of-the-art LLMs generally struggle to synthesize correct or sufficiently strong invariants to accelerate program verification, the strongest model o3 can still yield non-trivial improvements on a meaningful fraction of problems.
-Table˜3 reports results on the hard split of InvBench.
-In this setting, LLMs show only negligible improvement over UAutomizer.
-The only noteworthy exception is gpt-oss-120b, which produces an invariant for one problem that delivers a 29.57× speedup.
-However, overall, such speedup cases are exceedingly rare.
-There are three main takeaways from the results.
-First, generating strong invariants that yield performance speedups is substantially more difficult than merely producing correct invariants, as reflected in the large gap between the percentage of correct invariants and the percentage of speedups.
-Second, model capability is a key factor, as demonstrated by the sharp differences in speedups reported in Table˜2.
-Third, LLMs remain far from fully addressing the task of invariant synthesis, leaving considerable room for future progress, as shown in Table˜3.
-5.2 Results of Verifiers on InvBench
-We compare the state-of-the-art non-LLM-based tool UAutomizer (Schüssele et al., 2024) with other LLM-based verifiers on both the easy split and the hard split of InvBench.
-On the easy split of InvBench, where UAutomizer solves all 113 instances within 30 seconds, it clearly surpasses all LLM-based verifiers.
-LaM4Inv (Wu et al., 2024a) and Loopy (Kamath et al., 2023) fail to solve any instance within this time limit, while LEMUR (Wu et al., 2024b) solves only 55 instances.
-These results indicate that none of the LLM-based verifiers provide any advantage over UAutomizer on the easy split.
-Table˜4 presents the comparison on the hard split of InvBench.
-UAutomizer still delivers the best performance across almost all timeouts.
-UAutomizer consistently and significantly outperforms LaM4Inv and Loopy.
-In particular, LaM4Inv and Loopy solve fewer than half as many instances as UAutomizer across all timeouts.
-Notably, within 10 seconds, LEMUR solves more problems than UAutomizer, suggesting that it can accelerate some of the instances.
-Nevertheless, UAutomizer still outperforms LEMUR with longer timeouts, indicating that although LEMUR can offer early solving advantages on certain challenging instances, its overall capability is still limited compared to UAutomizer.
-Given the consistently poor performance of LaM4Inv, which often either fails or times out, we conducted a manual investigation.
-Our analysis shows that LaM4Inv fails on some examples when applied to external datasets beyond those used in its customized evaluation.
-However, the preprocessing steps or manual annotations required by the tool are not documented.
-The official repository does not provide preprocessing scripts, and our attempt to contact the authors has not received a response.
-For a direct comparison between UAutomizer and LaM4Inv, we also refer readers to Section˜5.3, where we evaluate UAutomizer on the dataset released by LaM4Inv.
-5.3 Results of UAutomizer on Prior Datasets
-As shown in Table 5, UAutomizer, the state-of-the-art non-LLM-based verifier, consistently solves more instances than LaM4Inv (Wu et al., 2024a) and Loopy (Kamath et al., 2023), two representative LLM-based tools, on their respective customized datasets.
-This highlights that prior work omitted an important baseline comparison against UAutomizer.
-LEMUR (Wu et al., 2024b) is the only tool that surpasses UAutomizer.
-However, this is largely explained by its dataset construction: LEMUR evaluates only on problems that UAutomizer cannot solve within 600 seconds.
-To assess whether this advantage generalizes beyond its own curated benchmark, we conduct an additional analysis presented in Section˜6.
-While LaM4Inv and Loopy consistently underperform UAutomizer in terms of the total number of solved instances across time budgets, they nevertheless offer complementary benefits, which we discuss in Section˜A.2.
-We further investigate the distributional differences between datasets.
-At the 600-second timeout, the performance of prior LLM-based verifiers approaches that of UAutomizer, with LaM4Inv even matching it exactly.
-In contrast, on InvBench (see Section˜5.2), all LLM-based verifiers perform poorly at 600 seconds, suggesting a substantial shift in distribution compared to the customized datasets used in prior work.
-Our analysis confirms this: programs in InvBench are significantly longer, averaging 62 lines of code in the hard split, compared to only 22 for LaM4Inv, 23 for LEMUR, and 27 for Loopy.
-Manual inspection shows that InvBench contains features such as multiple loops, functions, arrays, and pointers, which are largely absent from prior datasets.
-This makes InvBench a more challenging and realistic benchmark that better distinguishes solver performance.
-We also note that different LLM-based verification frameworks are built on different base solvers: LEMUR is built on UAutomizer, Loopy on Frama-C (Cuoq et al., 2012), and LaM4Inv on ESBMC (Gadelha et al., 2018).
-Given the strength of UAutomizer, we believe future work should place greater emphasis on developing LLM-based verifiers atop state-of-the-art solvers such as UAutomizer and ensure that comparisons against it are not omitted.
-Verifier Solved Instances under Different Timeouts
-10s 100s 300s 600s
-LaM4Inv (Wu et al., 2024a) 0 0 0 0
-Loopy (Kamath et al., 2023) 0 2 18 50
-LEMUR (Wu et al., 2024b) 32 71 73 73
-UAutomizer 0 73 95 113
-Table 4: Comparison of solved instances on InvBench-Hard by different verifiers under varying timeout budgets.
-Verifier Total Instances Solved Instances under Different Timeouts
-10s 100s 300s 600s
-LaM4Inv (Wu et al., 2024a) 316 144 286 295 299
-UAutomizer 316 299 299 299 299
-Loopy (Kamath et al., 2023) 469 0 133 353 403
-UAutomizer 469 372 403 411 413
-LEMUR (Wu et al., 2024b) 47 2 8 16 19
-UAutomizer 47 0 0 0 0
-Table 5: Comparison of prior LLM-based verifiers and UAutomizer on their own customized dataset under different timeout budgets.
-5.4 Performance Gains from Invariants Generated by UAutomizer
-To quantify the potential speedup achievable when providing correct and strong invariants to solvers, we conduct a study using UAutomizer.
-We extract the invariants identified by UAutomizer and then measure the resulting speedup when they are supplied to the verifier.
-On a random sample of 100 problems from our training set, we observe an overall average speedup of 1.86×.
-This indicates that the invariants discovered by UAutomizer are sufficiently strong to accelerate verification, and we leverage them for fine-tuning LLMs in Section˜5.5.
-5.5 Effectiveness of Fine-Tuning
-For each synthetically generated problem in the training set, we run UAutomizer to produce invariants, yielding a training dataset of 3589 problems with associated invariants.
-We then perform supervised fine-tuning using LoRA for 3 epochs.
-[…]
-Qwen2.5-72B (base) 4.4% 0.9% 1.20× 1.00×
-Qwen2.5-72B (fine-tuned) 32.7% 26.5% 1.13× 1.03×
-Qwen3-Coder-480B (base) 14.2% 8.0% 1.09× 1.01×
-Qwen3-Coder-480B (fine-tuned) 40.7% 29.2% 1.29× 1.08×
-Table 6: Improvement of supervised fine-tuning on InvBench-Easy.
-Table˜6 reports results on the easy split of InvBench.
-Fine-tuning leads to substantial gains in both invariant correctness and runtime speedup.
-Qwen3-Coder-480B shows the most pronounced improvements: the proportion of correct invariants increases from 14.2% in the base model to 40.7%, and the percentage of instances with speedups rises from 8% to 29.2%.
-For Qwen2.5-72B, the conditional speedup decreases after fine-tuning, but this is because the baseline conditional speedup was driven by a single case, whereas fine-tuning yields a larger number of cases with speedups.
-We also note that on the hard split, neither of the fine-tuned models shows improvement.
-These results suggest that more training efforts are required to achieve gains on harder instances.
-5.6 Effectiveness of Best-of-N Sampling
-[…]
-claude-opus-4.1 1 18.6% 8.8% 1.23× 1.02×
-claude-opus-4.1 16 37.2% 23.0% 1.15× 1.03×
-claude-sonnet-4 1 15.0% 8.8% 1.06× 1.01×
-claude-sonnet-4 16 36.3% 22.1% 1.20× 1.04×
-o3 1 39.8% 28.3% 1.37× 1.09×
-o3 16 50.4% 35.4% 1.39× 1.12×
-Table 7: Results of Best-of-N sampling (N = 16) on InvBench-Easy.
-We also evaluated whether repeated sampling improves LLM performance in invariant synthesis.
-For each problem, we generated 16 samples at a temperature of 0.7, removed duplicate invariants, and verified all candidates in parallel.
-We then select the sample with the shortest verification time and kill all other samples.
-Table 7 summarizes the results on InvBench’s easy split, demonstrating that repeated sampling leads to consistent gains.
-For example, o3 improves from 39.8% to 50.4% correct invariants, while claude-opus-4.1 nearly doubles its accuracy from 18.6% to 37.2%.
-On the hard split, for claude-opus-4.1, with 16 samples, the percentage of instances with correct invariants increases from 14.2% to 15.9%, the percentage of instances with speedup from 0.9% to 2.7%, and the overall average speedup from 1.02× to 1.03×.
-Similarly, for claude-sonnet-4, the percentage of instances with correct invariants increases from 13.3% to 20.4%, the percentage of instances with speedup from 0.9% to 2.7%, conditional average speedup from 3.35× to 6.12×, and the overall average speedup from 1.01× to 1.07×.
-These improvements suggest that best-of-N sampling is a promising technique for LLMs to improve invariant synthesis performance.
-6 Discussion
-Generalizability of LEMUR.
-LEMUR (Wu et al., 2024b) reports the best results on its customized dataset, as it targets instances that UAutomizer cannot solve within 600 seconds.
-To test whether this advantage generalizes, we sampled 50 unsolved instances and found that LEMUR solved 12 within the same timeout of 600 seconds.
-This suggests that LEMUR’s gains are not solely due to benchmark design but reflect a genuine advantage on harder problems.
-Inference Overhead of Models.
-Our evaluation includes LLM serving time as a realistic measure of end-to-end performance.
-Since the goal is to accelerate verification, inference overhead must be considered alongside solver runtime.
-Future research may explore how to balance the quality of generated invariants with the inference cost of producing them.
-7 Conclusion
-This work introduced InvBench, a principled framework for evaluating the capabilities of LLMs in invariant synthesis.
-Our approach employs a verifier-based decision procedure with a formal soundness guarantee and assesses not only correctness but also the speedups that invariants contribute to program verification.
-Using a benchmark of 226 instances, we conducted a comparison across state-of-the-art LLMs, existing LLM-based verifiers, and the traditional solver UAutomizer.
-The results show that although LLM-based verifiers represent a promising direction, they do not yet offer significant advantages over non-LLM-based approaches.
-Model capability proves to be a critical factor, and our benchmark remains an open challenge for current LLMs.
-At the same time, we demonstrated that supervised fine-tuning and Best-of-N sampling can improve model performance in accelerating verification.
-LLM Usage
-LLMs are the subject of this study. We additionally used them for polishing the writing.
-Appendix A Appendix
-A.1 Proof
 Theorem (Decision Soundness).
@@
 ∎
-A.2 Complementary Results of LLM-based Solvers
-As shown in Table˜A1, all three LLM-based tools solve instances that remain unsolved by UAutomizer within the 600-second budget.
-LaM4Inv and Loopy each contribute additional solved cases, and LEMUR is able to handle 19 problems that UAutomizer cannot solve at all, underscoring the complementary strengths of LLM-based approaches.
-Verifier Solved Instances $\Delta$ Solved
-LaM4Inv 299 13
-Loopy 403 40
-LEMUR 19 19
-Table A1: LLM-based verifiers complement UAutomizer by solving problems beyond its reach.
-We report the total numbers of solved problems, and $\Delta$ Solved” is the number of instances uniquely solved that are unsolved by UAutomizer.
+This theorem establishes that whenever the calculus derives a conclusive outcome, that outcome is correct.
+The inconclusive case $\mathbf{U}$ is deliberately conservative: it makes no claim about the truth or falsity of the property and safely captures verifier incompleteness or timeouts.
+3.3 Implementation
+We describe the implementation of our verifier-based evaluation framework.
+Given a program $P$ and target property $p^{\star}$, the system evaluates the generated invariants according to the decision procedure.
+Each invariant $q=\langle\psi,\ell\rangle$ consists of a program location $\ell$ and predicate $\psi$.
+Syntactic Validation.
+Before invoking the verifier, we apply syntactic checks to the generated predicate $\psi$.
+These checks ensure that $\psi$ can be safely interpreted as a state predicate and that its insertion as an assumption does not alter the program state.
+For instance, expressions that update variables (e.g., a += 1) are rejected.
+Only Boolean conditions over the program state are accepted.
+Parallel Verifier Queries.
+For each candidate $q$, the procedure issues two verifier queries, namely $d_{a}=V(P,\varnothing,q)$ to check whether $q$ is an invariant and $d_{b}=V(P,\{q\},p^{\star})$ to check whether the target holds under the assumption $q$.
+These queries are executed in parallel in our implementation, which reduces latency and enables speedup when the proposed invariant is useful for verification.
+The final outcome is then derived exactly according to the decision calculus.
+3.4 Supervised Fine-Tuning and Best-of-N Sampling
+Below, we discuss how we construct our dataset for fine-tuning and the way we perform Best-of-N sampling.
+Synthetic Dataset Generation.
+To construct the synthetic dataset, we prompt GPT-4o using the template in Appendix A.1.
+The template takes three seed programs as examples and instructs the model to synthesize a new C program that is compilable and contains both loops and assertions.
+To obtain a diverse and large pool of candidates, we repeatedly invoke the model with different seed programs.
+To avoid data leakage, these seed programs are randomly drawn from the SV-COMP pool  (Beyer and Strejček, 2025) that is disjoint from our evaluation set.
+Although the prompt requests compilable programs with loops and assertions, the LLM-generated programs may fail to compile, include assertions that do not hold, and contain multiple assertions.
+For programs with multiple assertions, we split them into separate instances, each retaining only a single assertion while preserving all loop structures (ensuring at least one loop per instance).
+We then run UAutomizer on every program and discard any instance that is non-compilable or whose assertion is invalid.
+This filtering step ensures the quality of the dataset, resulting in 3589 synthetic programs.
+Extract Invariants Generated from UAutomizer.
+When running UAutomizer to prove the assertions in the synthetic programs, the tool also emits the invariants it discovers.
+From its output, we extract the loop invariants.
+Each extracted invariant includes its program location and its predicate.
+Although each program contains exactly one assertion, it may include multiple loops, so a single program can yield multiple loop invariants, all associated with the same assertion.
+We pair each program with each of its corresponding loop invariants to form our training dataset.
+We show an example invariant generated from UAutomizer in Appendix A.2.
+With the ground-truth invariants, we perform supervised fine-tuning using LoRA (Hu et al., 2022).
+Best-of-N Sampling.
+Best-of-N sampling is an inference-time strategy in which multiple candidate programs are generated, and the most effective one is selected, a technique shown to improve performance on code-generation tasks (Ehrlich et al., 2025).
+In our setting, the best candidate is the invariant that yields the largest speedup, i.e., the one whose decision procedure finishes earliest.
+As described in Section˜3.2, evaluating a single candidate requires two verifier queries issued in parallel; therefore, Best-of-N sampling evaluates 2N verifier queries concurrently.
+4 Experimental Setup
+Dataset.
+We construct our evaluation benchmark from SV-COMP (Beyer and Strejček, 2025), a standard competition in software verification, focusing on problems that require loop invariant synthesis.
+We collect a pool of 866 instances and run the state-of-the-art non-LLM verifier UAutomizer (Schüssele et al., 2024) with a 600-second timeout to record per-instance solving times.
+The process is fully automated, with no manual filtering or cherry-picking.
+We use this timeout to enable meaningful measurement of verification speedups.
+In contrast to prior work (Wu et al., 2024b), which evaluates only instances that the baseline cannot solve within this time budget, our approach requires knowing the baseline runtime to quantify relative improvements.
+Restricting to instances whose baseline runtime is observable allows us to report concrete speedup factors, rather than only counts of additionally solved problems.
+To the best of our knowledge, this dataset is the largest used in prior evaluations of LLM-based verifiers.
+Also, the programs in our dataset are significantly longer, with an average of 53 lines of code, compared to only 22 for LaM4Inv (Wu et al., 2024a), 23 for LEMUR (Wu et al., 2024b), and 27 for Loopy (Kamath et al., 2023).
+Manual inspection shows that our dataset contains features such as multiple loops, functions, arrays, and pointers, which are largely absent from prior datasets.
+This makes Quokka a more challenging and realistic benchmark that better distinguishes solver performance.
+For training data, we construct our training data with synthetic data generation filtered by the solvers, as described in Section˜3.4.
+We make sure that the seed programs used for synthetic data generation are disjoint from our evaluation dataset, mitigating the leakage issue.
+The statistics of our dataset are shown in Table˜1.
+Dataset Avg. #Lines #Instances
+Evaluation 53 866
+Training 42 3589
+Table 1: Dataset statistics.
+Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
+Qwen2.5-7B 419 34 26 23 20 18
+o3 550 38 33 30 24 23
+Llama-3.1-8B 342 43 35 32 29 27
+claude-opus-4.1 487 56 45 38 35 31
+Qwen2.5-72B 500 63 49 45 37 33
+claude-opus-4.5 689 69 53 46 39 34
+gpt-5.1 694 56 46 41 40 37
+claude-sonnet-4 620 68 58 51 45 41
+gpt-5.2 710 81 65 60 56 51
+Table 2: Performance of different LLMs when using Quokka.
+We report the number of instances with verified-correct invariants and the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over UAutomizer.
+Metrics.
+We evaluate LLMs along two dimensions: (1) the correctness of the generated invariants and (2) the performance improvements they provide.
+Correctness is determined by the solver using a timeout set to $1.2\times$ the original solving time for each problem.
+This slack allows the solver additional time to assess the correctness of model-generated invariants, reducing inconclusive outcomes caused by timeouts.
+All speedup-related measurements are reported relative to the original solver, which serves as the baseline.
+We report speedup using the metric of $\mathtt{\#Fast}_{p}$, defined as the number of instances for which verification completes at least $p\times$ faster than the baseline, with $p\in\{1.2,1.4,\ldots\}$.
+Additionally, we also report the number of instances solved under varying timeout budgets.
+Choice of Solvers.
+We use the state-of-the-art non-LLM-based solver UAutomizer (Schüssele et al., 2024) as our base solver.
+Our methodology requires that the solver can verify programs without externally-provided invariants, enabling us to measure speedup relative to a functional baseline.
+While prior work has explored other solvers such as Frama-C (Cuoq et al., 2012) and ESBMC (Gadelha et al., 2018), we select UAutomizer for two key reasons.
+First, UAutomizer represents the current state-of-the-art in automated verification, outperforming ESBMC on standard benchmarks in SV-COMP.
+Second, unlike Frama-C, which requires externally-provided invariants and cannot proceed without them, UAutomizer has built-in invariant synthesis capabilities that allow it to verify programs autonomously.
+To demonstrate the generality of Quokka’s methodology across different solvers, we also present experiments with ESBMC in Appendix A.5, where we show that Quokka successfully accelerates ESBMC as well.
+Models.
+We benchmark Claude models from Anthropic, GPT models from OpenAI, and models from the Qwen and LLaMA families (Hui et al., 2024; Yang et al., 2025; Touvron et al., 2023).
+Hardware and OS.
+Experiments were conducted on a node with Intel E5-2640 v4 CPUs, 256G main memory, running Ubuntu 20.04 LTS.
+5 Results
+5.1 Results of LLMs
+Table˜2 presents the performance of different LLMs when using Quokka.
+The results are sorted by the number of instances achieving at least $2.0\times$ speedup (# Fast2.0), which ranges from 18 instances for Qwen2.5-7B to 51 instances for gpt-5.2.
+Overall, the table demonstrates substantial variation in both correctness and speedup performance across different models, with larger and more recent models generally achieving better results.
+Effectiveness of Quokka.
+Our approach demonstrates the speedup capabilities across all evaluated models.
+Even smaller models like Qwen2.5-7B achieve speedups on 34 instances at the $1.2\times$ threshold, while the best-performing model gpt-5.2 accelerates 81 instances at this level and achieves $2.0\times$ or greater speedup on 51 instances.
+These results demonstrate that Quokka can effectively leverage LLMs to generate invariants that meaningfully accelerate verification.
+The consistent speedup gains across different model families, from 7B parameter models to frontier models, and the progressive improvement with model scale underscore the effectiveness of our approach.
+Correctness vs. Performance Gap.
+A striking observation is the substantial gap between generating correct invariants and achieving meaningful speedups.
+For example, gpt-5.2 produces 710 correct invariants but achieves $2.0\times$ speedup on only 51 instances, while claude-opus-4.5 generates 689 correct invariants yet delivers $2.0\times$ speedup on merely 34 instances.
+This gap highlights that correctness alone is insufficient: invariants must also be strong enough to meaningfully accelerate verification, requiring not only logical correctness but also strategic insight into which invariants will most effectively guide the solver.
+Model Inference Time.
+We include model inference time in our evaluation to provide a fair and realistic assessment of end-to-end performance.
+As documented in Appendix A.4, smaller models such as Llama-3.1-8B and Qwen2.5-7B exhibit negligible generation times (0.3–0.4 seconds), while most models complete inference within 3 seconds.
+The o3 model has the highest overhead at 8.1 seconds on average.
+Importantly, unlike LOOPY (Kamath et al., 2023), which invokes the LLM up to 15 times per program and thus spends a substantial portion of its time budget on API calls, our approach requires only a single LLM invocation per program.
+Consequently, LLM inference accounts for only a small fraction of total runtime.
+Furthermore, our evaluation dataset contains sufficiently challenging programs (with solving times up to 600 seconds) such that solver time naturally dominates inference overhead, making the comparison meaningful even when inference latency is included.
+Prompting-Based Methods.
+We investigate the impact of different prompting strategies on invariant generation quality using gpt-5.2.
+As detailed in Appendix A.6, we evaluate 0-shot, 2-shot, 4-shot, and chain-of-thought (CoT) prompting.
+Few-shot prompting does not consistently improve performance, while CoT prompting yields the weakest results, likely due to increased generation time without proportional quality gains.
+These findings indicate that simple zero-shot prompting is most effective for invariant synthesis.
+5.2 Comparison with Prior LLM-based Verifiers
+Method # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
+LaM4Inv 1 1 1 1 1
+LOOPY 8 7 5 3 1
+LEMUR-original 19 16 13 12 10
+LEMUR-gpt-5.2 39 30 28 25 22
+Quokka-gpt-5.2 81 65 60 56 51
+Table 3: Comparison of speedup performance across different methods.
+We report the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over UAutomizer.
+Table˜3 compares the speedup performance of Quokka against prior LLM-based verifiers across different speedup thresholds, while Figure˜2 shows the number of solved instances over time for each method.
+Quokka Outperforms Prior LLM-based Verifiers.
+Quokka with gpt-5.2 achieves state-of-the-art results across all speedup metrics.
+Among prior work, LEMUR (Wu et al., 2024b) represents the strongest baseline.
+To ensure a fair comparison, we enhance LEMUR by replacing its original model (gpt-4) with gpt-5.2, creating LEMUR-gpt-5.2.
+Even with this enhancement, Quokka-gpt-5.2 significantly outperforms LEMUR-gpt-5.2: at the $1.2\times$ speedup threshold, Quokka-gpt-5.2 accelerates 81 instances versus 39 for LEMUR-gpt-5.2 and 19 for LEMUR-original.
+At $2.0\times$ speedup, Quokka-gpt-5.2 achieves 51 instances compared to 22 for LEMUR-gpt-5.2 and 10 for LEMUR-original.
+LaM4Inv (Wu et al., 2024a) and LOOPY (Kamath et al., 2023) demonstrate minimal speedup capabilities, achieving at most 1 and 8 instances respectively at the $1.2\times$ level.
+These results demonstrate that Quokka establishes a new state-of-the-art for LLM-based verification acceleration.
+Differences in Base Solvers.
+Different LLM-based verification frameworks are built on different base solvers: Quokka and LEMUR use UAutomizer, LOOPY uses Frama-C (Cuoq et al., 2012), and LaM4Inv uses ESBMC (Gadelha et al., 2018).
+The choice of base solver significantly impacts the performance, as stronger solvers provide a better foundation for speedup.
+Since UAutomizer represents the current state-of-the-art solver, we believe future work should prioritize developing LLM-based verifiers atop state-of-the-art solvers such as UAutomizer.
+For a direct comparison between UAutomizer and prior LLM-based solvers, we also refer readers to Appendix A.3, where we evaluate UAutomizer on the dataset released by each of the prior works.
+Algorithmic Simplicity and Parallelism.
+Prior work predominantly adopts complex, sequential, and iterative processes with highly customized algorithms to handle LLM-generated invariants.
+In contrast, Quokka employs a simple, verifier-based, and parallel approach that directly leverages the base solver’s capabilities.
+LaM4Inv uses a “query-filter-reassemble” strategy that filters and combines model-generated predicates through a complex algorithm, rather than using them directly.
+LOOPY employs a highly-customized Houdini algorithm with iterative fixing to filter candidates.
+LEMUR is closest to our approach, also directly using the base solver to verify the correctness of invariants, but follows a purely sequential decision process with backtracking.
+In contrast, Quokka directly uses LLM-generated invariants without decomposition or reassembly, and exploits parallelism by issuing the verification queries in parallel.
+This design reduces implementation complexity while enabling more efficient hardware utilization, contributing to the superior speedup performance demonstrated in our experiments.
+Figure 2: Number of instances solved by different methods over varying timeout budgets.
+Model # Incorrect Invariants # Goal Timeout # Invariant Timeout # Both Timeout
+gpt-5.1 78 96 245 110
+claude-sonnet-4 110 160 209 145
+gpt-5.2 56 139 230 121
+Table 4: Failure mode breakdown for the top three models on Quokka.
+Incorrect Invariants: the invariant is refuted by the solver.
+Goal Timeout: the invariant is verified, but proving the assertion under it times out.
+Invariant Timeout: verifying the invariant itself times out.
+Both Timeout: both verification queries time out.
+Model # Correct Invariants # Fast1.0 # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
+Qwen2.5-7B (base) 419 116 34 26 23 20 18
+Qwen2.5-7B (finetuned) 431 141 39 27 26 24 20
+Table 5: Performance comparison of base and fine-tuned Qwen2.5-7B. Fine-tuning leads to modest improvements in both correctness and speedup metrics.
+5.3 Failure Mode Analysis
+To understand why many invariants fail to accelerate verification in Quokka, we analyze failure modes for the top three models.
+We categorize failures into four types: 1) Incorrect Invariants: the invariant is refuted by the solver; 2) Goal Timeout: the invariant is verified, but proving the assertion under it times out; 3) Invariant Timeout: verifying the invariant itself times out; 4) Both Timeout: both verification queries time out.
+The results in Table˜4 reveal that while some failures stem from incorrect invariants, the dominant issue is timeout-related.
+This suggests that current models lack an understanding of what makes verification queries tractable for symbolic solvers.
+Future work should explore training reward models or other mechanisms to help models internalize solver difficulty, enabling them to propose invariants that genuinely simplify verification by decomposing hard problems into easier subproblems rather than inadvertently generating verification conditions that are as difficult or harder to solve than the original proof goal.
+Figure 3: Effect of Best-of-N sampling on the number of instances under different Fastp settings.
+5.4 Results of Fine-Tuning and Best-of-N Sampling
+Modest Improvement from Supervised Fine-Tuning.
+We perform supervised fine-tuning using LoRA (Hu et al., 2022) on Qwen2.5-7B for 3 epochs.
+As shown in Table˜5, fine-tuning yields modest improvements: correct invariants increase from 419 to 431, and instances achieving meaningful speedups improve across all thresholds.
+For example, Fast1.0 improves from 116 to 141, and Fast1.2 improves from 34 to 39.
+These results suggest that domain-specific fine-tuning can improve both correctness and speedup, though gains remain moderate.
+Benfits of Best-of-N Sampling.
+We evaluate Best-of-N sampling by generating multiple invariant candidates from gpt-5.2 at temperature 0.7 and selecting the one yielding the fastest verification time.
+As shown in Figure˜3, increasing N from 1 to 4 consistently improves speedup across all Fastp thresholds.
+However, performance plateaus beyond N=8 due to hardware constraints: our 20-core CPU must verify 2N queries in parallel, and when N exceeds 10, resource contention may negate sampling benefits.
+Memory limits can further constrain scalability.
+While Best-of-N sampling is effective, its benefits are hardware-bounded, though modest values of N yield meaningful improvements.
+6 Conclusion
+This work introduced Quokka, a simple yet effective framework for LLM-based invariant synthesis that accelerates program verification.
+Our approach employs a principled verification procedure that decomposes the original proof goal into two independent subproblems and launches verifier queries in parallel, providing sound evaluation while achieving state-of-the-art speedup results.
+Using a benchmark of 866 instances, we evaluated 9 state-of-the-art LLMs across multiple model families and compared against prior LLM-based verifiers.
+The results demonstrate that Quokka consistently outperforms all existing LLM-based verifiers, achieving speedups of at least 1.2× on 81 instances compared to 39 instances for the previous best approach.
+We further demonstrated that supervised fine-tuning and Best-of-N sampling can yield measurable improvements in accelerating verification.
+Our work establishes a new state-of-the-art for LLM-based program verification and provides a principled foundation for future research in this direction.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix A Appendix
+A.1 Prompt Template Used for Training Dataset Generation
+Table˜A1 shows the prompt template used for synthesizing training programs from seed programs.
+You will be shown 3 example C programs.
+Please gain inspiration from the following programs to create a new high-quality C program.
+Do not simply copy from any of them.
+Requirements for the generated program: 1.
+The program MUST contain non-trivial loops (for or while).
+2.
+The program MUST contain assertions.
+3.
+The program MUST be compilable, self-contained, and reasonably complex (not trivial or overly short).
+4.
+Only output the new C program.
+Example snippets: Program 1: {SEED_PROGRAM_1}
+Program 2:
+{SEED_PROGRAM_2}
+Program 3:
+{SEED_PROGRAM_3}
+Output format: The generated program must be wrapped strictly in the following format: ‘‘‘c <NEW_C_PROGRAM>
+Table A1: Prompt template for synthetic data generation.
+A.2 An Example from the Fine-Tuning Dataset
+Original Program (Input to UAutomizer)
+1int main() {
+2 int N = __VERIFIER_nondet_int();
+3 assume_abort_if_not(N >= 1 && N < 100);
+4 int x = 0, y = 0;
+5 for (int i = 0; i < N; i++) {
+6 if (i % 2 == 0) {
+7 x += i;
+8 } else {
+9 y += i;
+10 }
+11 }
+12 int diff = x - y;
+13 __VERIFIER_assert(diff <= N);
+14 return 0;
+15}
+Loop Invariant Generated by UAutomizer:
+1Line Number: 6
+2Predicate:
+3(
+4(i + 1) % 2 == 0 &&
+5x < 2 + i + y &&
+6x < 2 + y + 2 * i &&
+7x < N + y + 1 &&
+81 <= N
+9)
+10||
+11(
+12i % 2 == 0 &&
+13x < 2 + i + y &&
+14x < 2 + y &&
+151 <= N
+16)
+Figure A1: An example from the fine-tuning dataset: program and its loop invariant generated by UAutomizer.
+Figure˜A1 shows an example from the fine-tuning dataset with the program to the UAutomizer and the generated loop invariant.
+The loop invariant holds at Line 6 (the beginning of the loop).
+It is a disjunction of two clauses, and can be written as $I\equiv P\lor Q=\big((i+1)\bmod 2=0\land x<2+i+y\land x<2+y+2i\land x<N+y+1\land 1\leq N\big)\lor\big(i\bmod 2=0\land x<2+i+y\land x<2+y\land 1\leq N\big)$.
+By inspecting the loop, we can derive an exact relationship between $x$, $y$, and $i$ at the beginning of each iteration.
+Since $x$ accumulates all even numbers less than $i$ and $y$ accumulates all odd numbers less than $i$, we obtain:
+$\text{if $i$ is even:}\quad x-y=-\frac{i}{2},\qquad\text{if $i$ is odd:}\quad x-y=\frac{i-1}{2}.$
+Using this relationship, it is straightforward to verify that the invariant $I$ produced by UAutomizer is correct. Moreover, $I$ is strong enough to prove the final assertion $x-y\leq N$.
+To show that the assertion holds at loop termination, we can check the following verification condition
+$I\;\wedge\;(i=N)\;\Rightarrow\;(x-y\leq N),$
+where $i=N$ denotes the loop’s exit condition.
+Case 1: $N$ is odd.
+Instantiating the invariant with $i=N$ activates the $P$ disjunct of $I$, from which we obtain
+xy<N+1.x-y<N+1.
+Since xyx-y is an integer, this directly implies $x-y\leq N$.
+Case 2: $N$ is even.
+In this case, the $Q$ disjunct applies. From the clause $x<2+y$ contained in $Q$, we derive
+xy<2.x-y<2.
+Because xyx-y is an integer and $N\geq 1$ and even (hence $N\geq 2$), we conclude
+$x-y\leq 1\leq N,$
+establishing the desired post-condition.
+Thus, the invariant $I$ indeed suffices to prove the final assertion.
+A.3 Results of UAutomizer on Prior Datasets
+Verifier Total Instances Solved Instances under Different Timeouts
+10s 100s 300s 600s
+LaM4Inv (Wu et al., 2024a) 316 144 286 295 299
+UAutomizer 316 299 299 299 299
+Loopy (Kamath et al., 2023) 469 0 133 353 403
+UAutomizer 469 372 403 411 413
+LEMUR (Wu et al., 2024b) 47 2 8 16 19
+UAutomizer 47 0 0 0 0
+Table A2: Comparison of prior LLM-based verifiers and UAutomizer on their own custom dataset under different timeout budgets.
+As shown in Table A2, UAutomizer, the state-of-the-art non-LLM-based verifier, consistently solves more instances than LaM4Inv (Wu et al., 2024a) and Loopy (Kamath et al., 2023), two representative LLM-based tools, on their respective custom datasets.
+This demonstrates the importance of base solver choice.
+LEMUR (Wu et al., 2024b) is the only tool that surpasses UAutomizer on its own benchmark.
+This outcome is largely explained by LEMUR’s dataset construction methodology: it specifically curates problems that UAutomizer cannot solve within 600 seconds.
+To assess whether LEMUR’s advantage generalizes beyond this carefully selected benchmark, we conduct an additional analysis.
+We randomly selected 50 instances that UAutomizer failed to solve within 600 seconds and evaluated LEMUR’s performance on them.
+LEMUR successfully solved 12 of these 50 instances within the same 600-second timeout, demonstrating that its gains are not merely an artifact of benchmark design but reflect genuine capability on challenging problems.
+Nevertheless, for our work, we focus on problems that UAutomizer can solve within the time budget, as our primary objective is to quantify speedup.
+A.4 Inference Time of Different Models
+Table˜A3 presents the average inference time per instance for different LLMs when generating invariants.
+As expected, smaller open-source models such as Llama-3.1-8B and Qwen2.5-7B exhibit the fastest generation times, averaging around 0.3–0.4 seconds per instance.
+Mid-sized models like Qwen2.5-72B and the GPT-5 series require approximately 0.6–1.4 seconds.
+The Claude models take longer at around 1.6–2.7 seconds per instance.
+The o3 model has the highest inference overhead, averaging over 8 seconds per instance.
+These results highlight the trade-off between model capability and inference cost, which must be considered when evaluating end-to-end verification performance.
+Model Avg Generation Time (s)
+Llama-3.1-8B 0.300
+Qwen2.5-7B 0.355
+Qwen2.5-72B 0.578
+gpt-5.2 1.076
+gpt-5.1 1.373
+claude-opus-4.1 1.626
+claude-opus-4.5 2.538
+claude-sonnet-4 2.746
+o3 8.075
+Table A3: Average inference time per instance for different LLMs when generating invariants.
+Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
+gpt-5.1 399 10 9 9 9 9
+gpt-5.2 455 17 11 10 9 9
+Table A4: Performance of different LLMs when using ESBMC as the base verifier.
+We report the number of instances with verified-correct invariants and the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over ESBMC.
+A.5 Results for ESBMC as the Base Solver
+To demonstrate the generality of Quokka’s methodology, we conduct additional experiments using ESBMC (Gadelha et al., 2018) as the base solver instead of UAutomizer.
+As shown in Table˜A4, Quokka successfully accelerates ESBMC as well, with both gpt-5.1 and gpt-5.2 generating hundreds of correct invariants and achieving speedups on multiple instances.
+For example, gpt-5.2 produces 455 correct invariants and achieves speedups of at least $1.2\times$ on 17 instances.
+These results confirm that Quokka is not limited to a specific solver and can be applied to accelerate different verification tools.
+A.6 Results for Prompting-Based Methods
+Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
+gpt-5.2 (CoT) 639 41 32 26 22 20
+gpt-5.2 (4-shot) 756 72 62 53 47 43
+gpt-5.2 (2-shot) 764 83 67 58 53 49
+gpt-5.2 710 81 65 60 56 51
+Table A5: Comparison of different prompting strategies for gpt-5.2.
+We investigate the impact of different prompting strategies on invariant generation quality using gpt-5.2.
+As shown in Table˜A5, we compare chain-of-thought (CoT) prompting, few-shot prompting with 2 and 4 examples, and our default zero-shot approach.
+The results show that few-shot prompting does not necessarily improve performance and may introduce noise or distract the model from the task at hand.
+Chain-of-thought prompting yields the weakest results across all metrics, with only 639 correct invariants and 41 instances achieving $1.2\times$ speedup.
+This suggests that explicit reasoning steps may increase generation time without proportional benefits to invariant quality, thereby reducing overall speedup.
+These findings indicate that prompting-based strategy can hardly improve the task of invariant synthesis.

### v3

@@
 We investigate whether large language models (LLMs) can accelerate program verification by generating useful loop invariants.
-We introduce Quokka, a first-order and effective framework for LLM-based invariant synthesis that provides sound evaluation while achieving state-of-the-art speedup results.
-Unlike prior work that designs complex, highly customized algorithms, Quokka employs a simple and principled verification procedure.
-We construct a benchmark of 866 instances and evaluate 9 state-of-the-art LLMs across multiple model families.
-Our results show that Quokka consistently outperforms all prior LLM-based verifiers: achieving speedups of at least 1.2× on 81 instances compared to 39 instances for the previous best approach.
-We further demonstrate that supervised fine-tuning and Best-of-N sampling can yield measurable improvements in accelerating verification.
-Machine Learning, ICML
+We introduce Quokka, an evaluation-oriented framework for LLM-based invariant synthesis that provides sound evaluation and achieves state-of-the-art performance.
+Unlike prior work that treats LLM outputs as noisy symbolic material requiring substantial post-processing, Quokka adopts a simpler and evaluation-centric design that directly validates whether each LLM-generated invariant helps prove the target assertion.
+We construct a benchmark of 866 instances derived from SV-COMP and evaluate 9 state-of-the-art LLMs across multiple model families.
+We demonstrate that supervised fine-tuning and Best-of-N sampling yield measurable improvements, and we show that Quokka consistently outperforms prior LLM-based verifiers.
+Our code and data are publicly available at https://github.com/Anjiang-Wei/Quokka
+footnotetext: Correspondence to: anjiang@cs.stanford.edu
 1 Introduction
@@
 However, only strong invariants can reduce verification effort and lead to a speedup.
-For example, in Figure˜1, the invariant $x>0$ is correct but not strong enough to prove the final assertion $x\neq 700$, whereas $x\equiv 3\pmod{7}$ is both correct and sufficiently strong.
+For example, in Figure 1, the invariant $x>0$ is correct but not strong enough to prove the final assertion $x\neq 700$, whereas $x\equiv 3\pmod{7}$ is both correct and sufficiently strong.
 Discovering such invariants is difficult and undecidable in general, which has motivated a long line of research.
@@
 Instead of evaluating LLMs in isolation, these efforts develop verification frameworks powered by LLMs.
-This line of work predominantly adopts complex, sequential, and iterative processes with highly customized algorithms to handle LLM-generated invariants.
+This line of work predominantly adopts complex algorithms to handle LLM-generated invariants, treating LLM output as noisy symbolic material that must be algorithmically repaired or reassembled.
 For instance, LaM4Inv (Wu et al., 2024a) uses a “query-filter-reassemble” strategy that filters and combines model-generated predicates through complex logical operations (conjunctions or disjunctions) to construct valid loop invariants, rather than using them directly.
+Clause2Inv (Cao et al., 2025) extends this idea further.
 LOOPY (Kamath et al., 2024) employs a customized Houdini algorithm with iterative refinement to filter candidate invariants.
-LEMUR (Wu et al., 2024b) develops a backtracking and sequential algorithm.
-While these approaches demonstrate the potential of LLM-based verification, their complexity raises a question: can we design a simpler, principled, and more effective decision procedure for LLM-based verifiers?
+LEMUR (Wu et al., 2024b) develops a recursive and backtracking algorithm for fixing invariants.
+While these approaches demonstrate the potential of LLM-based verification, their complexity raises a fundamental question: as models become increasingly capable, do we really need such elaborate, harness-heavy designs for building LLM-based verifiers?
+What if we stop rescuing LLM-generated invariants and instead measure directly whether a candidate invariant helps the solver?
 Figure 1: Illustration of Quokka’s evaluation pipeline.
 The LLM proposes an invariant by specifying a program location and predicate (e.g., location B with x % 7 == 3).
-The verification procedure then incorporates this invariant to prove the property x != 700 using two verifier queries, and we measure the resulting speedup relative to a baseline without LLM assistance.
-In this work, we introduce Quokka, a framework for LLM-based invariant synthesis that accelerates program verification.
-Quokka provides a sound approach to evaluate LLMs’ capabilities in generating invariants while achieving state-of-the-art speedup results compared with prior LLM-based verifiers.
-We formalize our methodology as a proof calculus, providing a rigorous foundation for building LLM-based verifiers.
-Instead of checking the correctness of invariants based on string-matching, we adopt a verifier-based approach that directly queries the underlying solver.
-Unlike previous work that designs complex, highly customized algorithms, Quokka employs a simple and principled verification procedure.
-Our approach enables both soundness guarantees and precise measurement of performance speedup, allowing us to systematically evaluate whether LLM-generated invariants can accelerate verification in practice.
+The verification algorithm then incorporates this invariant to prove the property x != 700 using two verifier queries, and we measure the resulting speedup relative to a baseline without LLM assistance.
+In this work, we introduce Quokka, an evaluation-oriented framework for LLM-based invariant synthesis that accelerates program verification.
+Prior LLM-based verification systems treat LLM outputs as imperfect intermediate artifacts requiring substantial symbolic post-processing—predicate filtering, clause combination, Houdini pruning, repair, or backtracking.
+In contrast, Quokka adopts a simpler, evaluation-centric design: it directly evaluates LLM-generated invariants by checking whether each invariant is valid and whether it helps prove the target assertion.
+This design eliminates most customized post-processing logic, enables parallel verification queries, and shifts the objective from merely finding correct invariants to finding invariants that can accelerate verification.
+Despite its simplicity, Quokka achieves state-of-the-art performance compared to all prior LLM-based verifiers.
 To support systematic comparison across solvers and LLMs, we construct a dataset of 866 instances derived from the most recent edition of the software verification competition SV-COMP (Beyer and Strejček, 2025).
-To the best of our knowledge, this is the largest evaluation dataset for LLM-based verifiers to date, compared with datasets used in previous work.
-Quokka consistently outperforms all prior LLM-based verifiers.
-When comparing against LEMUR (the previous best LLM-based verifier with the same underlying solver), Quokka achieves speedups of at least 1.2× (i.e., verification completes at least 20% faster than the baseline solver) on 81 instances compared to LEMUR’s 39 instances.
-For more substantial speedups of at least 2.0× (i.e., verification completes at least twice as fast), Quokka reaches 51 instances compared to LEMUR’s 22 instances.
-We evaluate 9 state-of-the-art LLMs, including GPT models from OpenAI, Claude models from Anthropic, and models from the Qwen and LLaMA families.
-Our experiments reveal that gpt-5.2 achieves the strongest speedup results among all evaluated models.
+To the best of our knowledge, this is the largest evaluation dataset for LLM-based verifiers to date.
+After evaluating 9 LLMs, we find that Quokka proposes a meaningful and challenging setup for LLMs, where model capabilities can be effectively distinguished and there remains significant room for improvement on challenging cases.
 Beyond model evaluation, we investigate techniques to enhance invariant generation capabilities.
@@
 Our experiments demonstrate that both supervised fine-tuning and Best-of-N sampling yield measurable improvements in accelerating verification.
-Specifically, Best-of-8 sampling achieves speedups of at least 1.2× on 99 instances, representing a 22% improvement over single-sample generation, which achieves such speedups on only 81 instances.
 In summary, our contributions are as follows:
-  • We introduce a decision procedure for LLM-based invariant synthesis that provides sound evaluation and achieves the best speedup results compared with prior LLM-based verifiers.
-  • We construct a benchmark of 866 instances, and conduct an evaluation of 9 state-of-the-art LLMs across multiple model families.
+  • We introduce a method for LLM-based invariant synthesis that provides sound evaluation and achieves the best performance results compared with prior LLM-based verifiers.
+  • We construct a benchmark of 866 instances, and conduct an evaluation of 9 LLMs across multiple model families.
   • We construct training data and demonstrate that both supervised fine-tuning and Best-of-N sampling yield measurable improvements in accelerating verification.
@@
 However, such invariants may fail to generalize to all possible executions, thereby compromising soundness.
-Our approach instead employs a verifier-based decision procedure relying on UAutomizer (Schüssele et al., 2024) that ensures soundness.
+Our approach instead employs a verifier-based algorithm relying on UAutomizer (Schüssele et al., 2024) that ensures soundness.
 Learning-Based Method for Invariant Generation.
-Machine learning based techniques have been widely adopted in invariant synthesis, including decision tree (Garg et al., 2014, 2016; Ezudheen et al., 2018; Riley and Fedyukovich, 2022; Xu et al., 2020), support vector machine (Li et al., 2017; Sharma et al., 2012), reinforcement learning (Si et al., 2018; Yu et al., 2023), and others (Sharma et al., 2013; Ryan et al., 2019; Yao et al., 2020).
-More recently, large language models have demonstrated strong capabilities in reasoning about code and logic (Wei et al., 2025a, b, c), giving rise to a series of work that explore using LLMs for finding invariants.
-(Pei et al., 2023) is the first pioneering work that evaluates LLMs’ capabilities in finding invariants, but it is not a sound evaluation.
+Machine learning based techniques have been widely adopted in invariant synthesis, including decision tree (Garg et al., 2014; 2016; Ezudheen et al., 2018; Riley and Fedyukovich, 2022; Xu et al., 2020), support vector machine (Li et al., 2017; Sharma et al., 2012), reinforcement learning (Si et al., 2018; Yu et al., 2023), and others (Sharma et al., 2013; Ryan et al., 2019; Yao et al., 2020).
+More recently, large language models have demonstrated strong capabilities in reasoning about code and logic (Wei et al., 2025a; b; c), giving rise to a series of work that explore using LLMs for finding invariants.
+Pei et al.
+(2023) is the first pioneering work that evaluates LLMs’ capabilities in finding invariants, but it is not a sound evaluation.
 Various techniques have been proposed to couple LLMs with symbolic solvers, including ranking LLM-generated invariants (Chakraborty et al., 2023), the “query-filter-reassemble” strategy of LaM4Inv (Wu et al., 2024a), the back-tracking algorithm in LEMUR (Wu et al., 2024b), and Loopy’s integration of the classic Houdini algorithm (Kamath et al., 2024).
-On the dataset side, (Liu et al., 2024) introduces a rule-based method for constructing a fine-tuning corpus, which differs from our verifier-based approach.
-In contrast, our work provides a simple and sound evaluation procedure for assessing LLM-generated invariants and investigates how both fine-tuning and Best-of-N sampling can enhance LLM performance in invariant synthesis.
+On the dataset side, Liu et al.
+(2024) introduces a rule-based method for constructing a fine-tuning corpus, which differs from our verifier-based approach.
+In contrast, our work provides a simple and sound evaluation algorithm for assessing LLM-generated invariants and investigates how both fine-tuning and Best-of-N sampling can enhance LLM performance in invariant synthesis.
 3 Method
@@
 Weak but correct invariants contribute little, leaving most of the reasoning to the verifier, whereas strong invariants narrow the search space of program states, reduce solver effort, and yield substantial speedups.
-3.2 Verifier-Based Decision Procedure
-We formalize our verifier-based procedure for assessing candidate invariants.
+3.2 Verifier-Based Approach
+We formalize our verifier-based algorithm for assessing candidate invariants.
 Let $P$ denote a program.
@@
 Given $P$ and $p^{\star}$, a large language model proposes a candidate invariant $q=\langle\psi,\ell\rangle$, typically at a loop header.
-To evaluate the utility of $q$, the procedure issues two verifier queries:
+To evaluate the utility of $q$, the algorithm issues two verifier queries:
 $d_{a}:=V(P,\varnothing,q)$
@@
 The query $d_{a}$ verifies whether $q$ is a correct invariant, while the query $d_{b}$ determines whether the target property holds under the assumption that $q$ is true.
-An illustration of these two verifier queries is provided in Figure˜1: $d_{a}$ corresponds to assert(x % 7 == 3) (upper subfigure), and $d_{b}$ corresponds to verifying the final assertion with the assume statement inserted (lower subfigure).
-The outcome of the procedure is expressed as a judgment
+An illustration of these two verifier queries is provided in Figure 1: $d_{a}$ corresponds to assert(x % 7 == 3) (upper subfigure), and $d_{b}$ corresponds to verifying the final assertion with the assume statement inserted (lower subfigure).
+The outcome of the algorithm is expressed as a judgment
 $P\;\Rightarrow\;\langle p^{\star},q\rangle\;\Downarrow\;d\qquad\text{with}\quad d\in\{\mathbf{T},\mathbf{F},\mathbf{U}\}.$
@@
 Rule DEC-$\mathbf{U}$ gives explicit conditions for inconclusiveness: the goal is not refuted under $q$ and $q$ is not established as an invariant.
+Theorem (Decision Soundness).
+If $P\Rightarrow\langle p^{\star},q\rangle\Downarrow\mathbf{T}$ is derivable, then $P\models p^{\star}$.
+If $P\Rightarrow\langle p^{\star},q\rangle\Downarrow\mathbf{F}$ is derivable, then $P\not\models p^{\star}$.
+We provide the proof in Appendix A.1.
+This theorem establishes that whenever the calculus derives a conclusive outcome, that outcome is correct.
+The inconclusive case $\mathbf{U}$ is deliberately conservative: it makes no claim about the truth or falsity of the property and safely captures verifier incompleteness or timeouts.
+3.3 Implementation
+We describe the implementation of our verifier-based evaluation framework.
+Given a program $P$ and target property $p^{\star}$, the system evaluates the generated invariants according to the approach.
+Each invariant $q=\langle\psi,\ell\rangle$ consists of a program location $\ell$ and predicate $\psi$.
+Syntactic Validation.
+Before invoking the verifier, we apply syntactic checks to the generated predicate $\psi$.
+These checks ensure that $\psi$ can be safely interpreted as a state predicate and that its insertion as an assumption does not alter the program state.
+For instance, expressions that update variables (e.g., a += 1) are rejected.
+Only Boolean conditions over the program state are accepted.
+Parallel Verifier Queries.
+For each candidate $q$, the algorithm issues two verifier queries, namely $d_{a}=V(P,\varnothing,q)$ to check whether $q$ is an invariant and $d_{b}=V(P,\{q\},p^{\star})$ to check whether the target holds under the assumption $q$.
+These queries are executed in parallel in our implementation, which reduces latency and enables speedup when the proposed invariant is useful for verification.
+The final outcome is then derived exactly according to the calculus.
+3.4 Supervised Fine-Tuning and Best-of-N Sampling
+Synthetic Dataset Generation.
+We prompt GPT-4o (template in Appendix A.3) with three seed programs to synthesize new C programs containing loops and assertions.
+Seed programs are randomly drawn from SV-COMP (Beyer and Strejček, 2025), disjoint from our evaluation set to prevent data leakage.
+Generated programs may fail to compile, contain invalid assertions, or include multiple assertions.
+We split multi-assertion programs into separate instances, each with a single assertion and all loops.
+We then run UAutomizer on each program and discard non-compilable instances or those with invalid assertions, yielding 3589 synthetic programs.
+Extract Invariants Generated from UAutomizer.
+We extract loop invariants from UAutomizer’s output when proving assertions in the synthetic programs.
+Each invariant includes its program location and predicate.
+Since programs may contain multiple loops, we pair each program with each of its loop invariants to form our training dataset.
+An example is shown in Appendix A.5.
+We perform supervised fine-tuning using LoRA (Hu et al., 2022) on these ground-truth invariants.
+Best-of-N Sampling.
+Best-of-N sampling generates multiple candidates and selects the most effective one (Ehrlich et al., 2025).
+We choose the invariant yielding the largest speedup.
+Since each candidate requires two parallel verifier queries (Section 3.2), Best-of-N evaluates 2N queries concurrently.
+4 Experimental Setup
+Dataset.
+We construct our evaluation benchmark from SV-COMP (Beyer and Strejček, 2025), a standard software verification competition, focusing on problems requiring loop invariant synthesis.
+By filtering instances containing loop keywords (while or for), we obtain 866 instances.
+To the best of our knowledge, this is the largest dataset used in prior LLM-based verifier evaluations.
+As shown in Table 1, our programs are significantly longer than prior work.
+Manual inspection reveals our dataset contains multiple loops, functions, arrays, and pointers (features largely absent from prior datasets), making it a more challenging benchmark that better distinguishes solver performance.
+For training data, we construct 3589 instances with synthetic data generation filtered by the solvers, as described in Section 3.4.
+Dataset Avg. #Lines #Instances Multi-Loops Multi-Functions Arrays/Pointers
+LaM4Inv (Wu et al., 2024a) 18 316
+Clause2Inv (Cao et al., 2025) 18 366
+LEMUR (Wu et al., 2024b) 20 180
+Loopy (Kamath et al., 2024) 25 638
+Quokka (Ours) 52 866
+Table 1: Dataset statistics.
+Metrics.
+We evaluate LLMs on two dimensions: (1) correctness of generated invariants and (2) performance improvements.
+Correctness is verified by the solver with a timeout of $1.2\times$ the original solving time, providing slack to reduce inconclusive outcomes.
+We report the number of additionally solved instances (unsolvable by the base solver), total solved instances, and average solving time under varying timeout budgets.
+All measurements represent end-to-end execution time, including LLM inference overhead.
+Choice of Solvers.
+We use UAutomizer (Schüssele et al., 2024), the state-of-the-art non-LLM-based solver, as our base solver.
+Our methodology requires that the solver can verify programs without externally-provided invariants, enabling speedup measurement relative to a functional baseline.
+We select UAutomizer over alternatives like Frama-C (Cuoq et al., 2012) and ESBMC (Gadelha et al., 2018) because it represents the current state-of-the-art (outperforming ESBMC on SV-COMP benchmarks) and, unlike Frama-C, has built-in invariant synthesis capabilities for autonomous verification.
+To demonstrate generality, we also present ESBMC experiments in Appendix A.8.
+Additional details such as models and hardware setup are provided in Appendix A.2.
+5 Results
+5.1 Results of LLMs
+Model # Correct Invariants 60s 500s
+$\Delta$ # Solved $\overline{T}$ (s) $\Delta$ # Solved $\overline{T}$ (s)
+Llama-3.1-8B 342 5 321 41.9 0 342 309.9
+Qwen2.5-7B 419 5 402 37.6 0 417 267.5
+claude-opus-4.1 487 13 466 34.2 1 487 229.4
+Qwen2.5-72B 500 11 474 33.8 0 500 221.7
+o3 550 17 534 35.9 1 549 198.3
+claude-sonnet-4 620 16 591 27.8 1 620 156.0
+claude-opus-4.5 689 17 659 24.2 1 689 118.8
+gpt-5.1 694 15 661 23.3 1 694 113.8
+gpt-5.2 710 21 681 22.2 1 710 105.1
+Table 2: Performance comparison of LLMs on invariant generation: number of correct invariants, additional instances solved beyond UAutomizer baseline ($\Delta$), total solved instances, and average solving time ($\overline{T}$) under 60s and 500s timeouts.
+Model # Correct Invariants 60s 500s
+# $\Delta$ # Solved $\overline{T}$ # $\Delta$ # Solved $\overline{T}$
+Qwen2.5-7B (base) 419 5 402 37.6 0 417 267.5
+Qwen2.5-7B (finetuned) 431 7 414 36.8 0 429 260.8
+Table 3: Performance comparison of base and fine-tuned Qwen2.5-7B. Fine-tuning leads to modest improvements in both correctness and speedup metrics.
+Table 2 presents the performance of different LLMs when using Quokka.
+Overall, the table demonstrates substantial variation in both correctness and performance across different models, with larger and more recent models generally achieving better results.
+Quokka proposes a meaningful and challenging problem setup for LLMs.
+The results in Table 2 reveal an interesting performance pattern across timeout budgets.
+Under the 60s timeout, LLMs solve much more instances beyond the UAutomizer baseline (ranging from 5 additional instances for smaller models to 21 for gpt-5.2), demonstrating their ability to accelerate verification on moderately difficult problems.
+However, under the 500s timeout, the incremental benefit diminishes significantly, indicating that LLMs struggle to provide meaningful invariants for the hardest problems that require more solving time.
+This pattern suggests that Quokka establishes a well-calibrated benchmark: LLMs already provide meaningful acceleration on a portion of instances, yet significant room remains for improvement on challenging cases.
+The substantial variation in correct invariants (342 for Llama-3.1-8B to 710 for gpt-5.2) and average solving times (41.9s to 22.2s under 60s timeout) further demonstrates that the benchmark effectively distinguishes model capabilities.
+While alternative verification algorithms might yield different absolute performance numbers, the superiority and simplicity of Quokka over other LLM-based verifiers (discussed in Section 5.4) establishes it as an appropriate framework for evaluating LLMs’ invariant generation capabilities.
+We provide further analysis of model inference time in Appendix A.7 and evaluate alternative prompting strategies (few-shot and chain-of-thought prompting) in Appendix A.9.
+5.2 Failure Mode Analysis
+To understand why many invariants fail to accelerate verification in Quokka, we analyze the failure modes of the top three models, with detailed results provided in Appendix A.11.
+The analysis shows that although some failures are caused by incorrect invariants, the majority are due to timeouts.
+We randomly sampled 50 timeout cases and manually analyzed the invariants.
+Our analysis reveals several key patterns.
+First, only 7 cases (14%) involved genuinely incorrect invariants, indicating reasonable LLM correctness.
+Second, 19 cases (38%) generated invariants identical to the original assertion, providing no decomposition benefit, with 4 additional cases nearly identical.
+Third, 9 cases (18%) generated trivial invariants (e.g., simple bounds) that failed to capture essential loop behavior.
+Fourth, only 5 cases (10%) produced strong, non-trivial invariants that still timed out, suggesting inherently difficult verification queries.
+Finally, 6 cases are incorrect original assertions where the LLM generated correct invariants.
+We provide case studies in Appendix A.12.
+5.3 Results of Fine-Tuning and Best-of-N Sampling
+Figure 2: Effect of Best-of-N sampling on the number of $\Delta$ instances solved over Uautomizer under different timeout settings.
+Modest Improvement from Supervised Fine-Tuning.
+We perform supervised fine-tuning using LoRA (Hu et al., 2022) on Qwen2.5-7B for 3 epochs.
+As shown in Table 3, fine-tuning yields modest improvements: correct invariants increase from 419 to 431, additionally solved instances increase from 5 to 7 under 60s timeout.
+These results suggest that domain-specific fine-tuning can improve both correctness and speedup, though gains remain moderate.
+Benefits of Best-of-N Sampling.
+We evaluate Best-of-N sampling by generating multiple invariant candidates from gpt-5.2 at temperature 0.7 and selecting the one yielding the fastest verification time.
+As shown in Figure 2, increasing N from 1 to 4 generally improves performance.
+The benefits continue to grow up to N=8, which achieves the best results.
+Beyond N=8, performance plateaus as resource contention and memory constraints begin to offset the advantages of additional sampling.
+These results demonstrate that Best-of-N sampling is an effective strategy for improving verification performance, with optimal gains observed at N=8 in our experimental setup.
+5.4 Comparison with Prior LLM-based Verifiers
+Method 30s 60s 250s 350s 500s
+# $\Delta$ $\overline{T}$(s) # $\Delta$ $\overline{T}$(s) # $\Delta$ $\overline{T}$(s) # $\Delta$ $\overline{T}$(s) # $\Delta$ $\overline{T}$(s)
+LaM4Inv 1 30.0 1 59.9 1 249.3 0 348.9 0 498.4
+Clause2Inv-original 14 25.8 9 49.9 4 199.2 2 277.4 0 394.4
+Clause2Inv-gpt-5.2 20 25.1 13 48.2 6 190.7 3 265.2 1 377.0
+Loopy 0 30.0 0 60.0 5 240.0 3 299.5 2 360.4
+LEMUR-original 7 28.6 7 50.4 3 165.3 1 222.3 1 306.8
+LEMUR-gpt-5.2 15 27.9 13 47.9 7 153.9 4 205.9 2 282.8
+Quokka-gpt-5.2 (best-of-8) 41 17.5 25 26.6 7 78.5 4 104.8 3 143.8
+Quokka-gpt-5.2 (2-shot) 33 14.3 20 19.4 6 45.2 2 57.6 2 76.0
+Table 4: Extra solved instances over UAutomizer and average solving time at different timeout thresholds.
+Table 4 compares Quokka against prior LLM-based verifiers, reporting additional instances solved beyond UAutomizer and average solving time under different timeout budgets.
+Figure 3 shows cumulative solved instances over time, demonstrating Quokka’s superior performance across all timeout thresholds.
+Method 30s 60s 250s 350s 500s
+# Solved $\overline{T}$(s) # Solved $\overline{T}$(s) # Solved $\overline{T}$(s) # Solved $\overline{T}$(s) # Solved $\overline{T}$(s)
+LEMUR-gpt-5.2 4 27.9 9 34.0 18 75.9 19 88.1 20 106.6
+Quokka-gpt-5.2 19 9.7 20 11.6 24 28.8 24 28.8 24 28.8
+Table 5: Performance comparison on the benchmark used in LEMUR.
+Since UAutomizer solves 0 instances within ten minutes on this dataset, # Solved equals # $\Delta$ (additional instances solved beyond the baseline).
+Results shown for different timeout thresholds.
+Quokka Outperforms Prior LLM-based Verifiers.
+Quokka with gpt-5.2 and best-of-8 sampling achieves state-of-the-art results across all speedup metrics, while Quokka with gpt-5.2 and 2-shot prompting also achieves leading performance under most timeout thresholds.
+Among prior work, LEMUR (Wu et al., 2024b) and Clause2Inv (Cao et al., 2025) represent the strongest baselines.
+LEMUR supports both UAutomizer and ESBMC.
+We evaluate LEMUR with UAutomizer, which is the state-of-the-art solver and matches the base solver used in our method.
+For fair comparison, we upgrade both LEMUR and Clause2Inv to gpt-5.2 (from gpt-4 and gpt-4o-mini, respectively), creating LEMUR-gpt-5.2 and Clause2Inv-gpt-5.2.
+Quokka-gpt-5.2 outperforms both upgraded baselines.
+LaM4Inv (Wu et al., 2024a) [^1] LaM4Inv’s performance differs from its original paper due to dataset differences and its requirement for manual preprocessing (confirmed by the authors), for which no automated script exists.
+and LOOPY (Kamath et al., 2024) show minimal gains over UAutomizer.
+Figure 3: Number of instances solved by different methods over varying timeout budgets.
+To further quantify the performance gap at larger timeout, we evaluate on LEMUR’s benchmark (Table 5)—challenging instances where UAutomizer fails within 10 minutes.
+Quokka consistently outperforms LEMUR-gpt-5.2 across all timeouts, solving substantially more instances with lower average solving times.
+Analysis.
+The baselines differ in both base solver choices and algorithm design.
+We analyze base verifier differences in Appendix A.10.
+Regarding algorithm design, Quokka adopts a distinctly “evaluation-oriented” approach: rather than implementing a handcrafted invariant search algorithm, it provides a clean interface between an LLM and a verifier.
+This represents a conceptual shift from prior work, which introduces complex algorithms to transform imperfect LLM outputs into usable invariants.
+Prior approaches treat LLM output as noisy symbolic material requiring algorithmic repair, whereas Quokka treats it as a direct proof candidate to be validated by the verifier.
+For instance, LaM4Inv requests complete invariants but then filters predicates with model checker and reassembles them, indicating distrust in the raw output.
+Clause2Inv takes this further.
+However, as models become increasingly capable, simpler evaluation-centric designs become preferable.
+We additionally provide a detailed comparison between Quokka and LEMUR in Appendix A.10.
+6 Conclusion
+This work introduced Quokka, an evaluation-oriented framework for LLM-based invariant synthesis that accelerates program verification.
+Unlike prior work that treats LLM outputs as noisy symbolic material requiring substantial post-processing, Quokka adopts a simpler and evaluation-centric design that directly validates whether each LLM-generated invariant helps prove the target assertion.
+We constructed a benchmark of 866 instances derived from SV-COMP and evaluated 9 state-of-the-art LLMs across multiple model families.
+Our results demonstrate that Quokka consistently outperforms prior LLM-based verifiers.
+We further showed that supervised fine-tuning and Best-of-N sampling yield measurable improvements in accelerating verification.
+Our work establishes a new state-of-the-art for LLM-based program verification and provides a principled foundation for future research in this direction.
+Appendix A Appendix
+A.1 Soundness Proof
 Theorem (Decision Soundness).
@@
 ∎
-This theorem establishes that whenever the calculus derives a conclusive outcome, that outcome is correct.
-The inconclusive case $\mathbf{U}$ is deliberately conservative: it makes no claim about the truth or falsity of the property and safely captures verifier incompleteness or timeouts.
-3.3 Implementation
-We describe the implementation of our verifier-based evaluation framework.
-Given a program $P$ and target property $p^{\star}$, the system evaluates the generated invariants according to the decision procedure.
-Each invariant $q=\langle\psi,\ell\rangle$ consists of a program location $\ell$ and predicate $\psi$.
-Syntactic Validation.
-Before invoking the verifier, we apply syntactic checks to the generated predicate $\psi$.
-These checks ensure that $\psi$ can be safely interpreted as a state predicate and that its insertion as an assumption does not alter the program state.
-For instance, expressions that update variables (e.g., a += 1) are rejected.
-Only Boolean conditions over the program state are accepted.
-Parallel Verifier Queries.
-For each candidate $q$, the procedure issues two verifier queries, namely $d_{a}=V(P,\varnothing,q)$ to check whether $q$ is an invariant and $d_{b}=V(P,\{q\},p^{\star})$ to check whether the target holds under the assumption $q$.
-These queries are executed in parallel in our implementation, which reduces latency and enables speedup when the proposed invariant is useful for verification.
-The final outcome is then derived exactly according to the decision calculus.
-3.4 Supervised Fine-Tuning and Best-of-N Sampling
-Below, we discuss how we construct our dataset for fine-tuning and the way we perform Best-of-N sampling.
-Synthetic Dataset Generation.
-To construct the synthetic dataset, we prompt GPT-4o using the template in Appendix A.1.
-The template takes three seed programs as examples and instructs the model to synthesize a new C program that is compilable and contains both loops and assertions.
-To obtain a diverse and large pool of candidates, we repeatedly invoke the model with different seed programs.
-To avoid data leakage, these seed programs are randomly drawn from the SV-COMP pool  (Beyer and Strejček, 2025) that is disjoint from our evaluation set.
-Although the prompt requests compilable programs with loops and assertions, the LLM-generated programs may fail to compile, include assertions that do not hold, and contain multiple assertions.
-For programs with multiple assertions, we split them into separate instances, each retaining only a single assertion while preserving all loop structures (ensuring at least one loop per instance).
-We then run UAutomizer on every program and discard any instance that is non-compilable or whose assertion is invalid.
-This filtering step ensures the quality of the dataset, resulting in 3589 synthetic programs.
-Extract Invariants Generated from UAutomizer.
-When running UAutomizer to prove the assertions in the synthetic programs, the tool also emits the invariants it discovers.
-From its output, we extract the loop invariants.
-Each extracted invariant includes its program location and its predicate.
-Although each program contains exactly one assertion, it may include multiple loops, so a single program can yield multiple loop invariants, all associated with the same assertion.
-We pair each program with each of its corresponding loop invariants to form our training dataset.
-We show an example invariant generated from UAutomizer in Appendix A.2.
-With the ground-truth invariants, we perform supervised fine-tuning using LoRA (Hu et al., 2022).
-Best-of-N Sampling.
-Best-of-N sampling is an inference-time strategy in which multiple candidate programs are generated, and the most effective one is selected, a technique shown to improve performance on code-generation tasks (Ehrlich et al., 2025).
-In our setting, the best candidate is the invariant that yields the largest speedup, i.e., the one whose decision procedure finishes earliest.
-As described in Section˜3.2, evaluating a single candidate requires two verifier queries issued in parallel; therefore, Best-of-N sampling evaluates 2N verifier queries concurrently.
-4 Experimental Setup
-Dataset.
-We construct our evaluation benchmark from SV-COMP (Beyer and Strejček, 2025), a standard competition in software verification, focusing on problems that require loop invariant synthesis.
-We collect a pool of 866 instances and run the state-of-the-art non-LLM verifier UAutomizer (Schüssele et al., 2024) with a 600-second timeout to record per-instance solving times.
-The process is fully automated, with no manual filtering or cherry-picking.
-We use this timeout to enable meaningful measurement of verification speedups.
-In contrast to prior work (Wu et al., 2024b), which evaluates only instances that the baseline cannot solve within this time budget, our approach requires knowing the baseline runtime to quantify relative improvements.
-Restricting to instances whose baseline runtime is observable allows us to report concrete speedup factors, rather than only counts of additionally solved problems.
-To the best of our knowledge, this dataset is the largest used in prior evaluations of LLM-based verifiers.
-Also, the programs in our dataset are significantly longer, with an average of 53 lines of code, compared to only 22 for LaM4Inv (Wu et al., 2024a), 23 for LEMUR (Wu et al., 2024b), and 27 for Loopy (Kamath et al., 2023).
-Manual inspection shows that our dataset contains features such as multiple loops, functions, arrays, and pointers, which are largely absent from prior datasets.
-This makes Quokka a more challenging and realistic benchmark that better distinguishes solver performance.
-For training data, we construct our training data with synthetic data generation filtered by the solvers, as described in Section˜3.4.
-We make sure that the seed programs used for synthetic data generation are disjoint from our evaluation dataset, mitigating the leakage issue.
-The statistics of our dataset are shown in Table˜1.
-Dataset Avg. #Lines #Instances
-Evaluation 53 866
-Training 42 3589
-Table 1: Dataset statistics.
-Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
-Qwen2.5-7B 419 34 26 23 20 18
-o3 550 38 33 30 24 23
-Llama-3.1-8B 342 43 35 32 29 27
-claude-opus-4.1 487 56 45 38 35 31
-Qwen2.5-72B 500 63 49 45 37 33
-claude-opus-4.5 689 69 53 46 39 34
-gpt-5.1 694 56 46 41 40 37
-claude-sonnet-4 620 68 58 51 45 41
-gpt-5.2 710 81 65 60 56 51
-Table 2: Performance of different LLMs when using Quokka.
-We report the number of instances with verified-correct invariants and the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over UAutomizer.
-Metrics.
-We evaluate LLMs along two dimensions: (1) the correctness of the generated invariants and (2) the performance improvements they provide.
-Correctness is determined by the solver using a timeout set to $1.2\times$ the original solving time for each problem.
-This slack allows the solver additional time to assess the correctness of model-generated invariants, reducing inconclusive outcomes caused by timeouts.
-All speedup-related measurements are reported relative to the original solver, which serves as the baseline.
-We report speedup using the metric of $\mathtt{\#Fast}_{p}$, defined as the number of instances for which verification completes at least $p\times$ faster than the baseline, with $p\in\{1.2,1.4,\ldots\}$.
-Additionally, we also report the number of instances solved under varying timeout budgets.
-Choice of Solvers.
-We use the state-of-the-art non-LLM-based solver UAutomizer (Schüssele et al., 2024) as our base solver.
-Our methodology requires that the solver can verify programs without externally-provided invariants, enabling us to measure speedup relative to a functional baseline.
-While prior work has explored other solvers such as Frama-C (Cuoq et al., 2012) and ESBMC (Gadelha et al., 2018), we select UAutomizer for two key reasons.
-First, UAutomizer represents the current state-of-the-art in automated verification, outperforming ESBMC on standard benchmarks in SV-COMP.
-Second, unlike Frama-C, which requires externally-provided invariants and cannot proceed without them, UAutomizer has built-in invariant synthesis capabilities that allow it to verify programs autonomously.
-To demonstrate the generality of Quokka’s methodology across different solvers, we also present experiments with ESBMC in Appendix A.5, where we show that Quokka successfully accelerates ESBMC as well.
+A.2 Additional Details on Experiment Setup
 Models.
@@
 Experiments were conducted on a node with Intel E5-2640 v4 CPUs, 256G main memory, running Ubuntu 20.04 LTS.
-5 Results
-5.1 Results of LLMs
-Table˜2 presents the performance of different LLMs when using Quokka.
-The results are sorted by the number of instances achieving at least $2.0\times$ speedup (# Fast2.0), which ranges from 18 instances for Qwen2.5-7B to 51 instances for gpt-5.2.
-Overall, the table demonstrates substantial variation in both correctness and speedup performance across different models, with larger and more recent models generally achieving better results.
-Effectiveness of Quokka.
-Our approach demonstrates the speedup capabilities across all evaluated models.
-Even smaller models like Qwen2.5-7B achieve speedups on 34 instances at the $1.2\times$ threshold, while the best-performing model gpt-5.2 accelerates 81 instances at this level and achieves $2.0\times$ or greater speedup on 51 instances.
-These results demonstrate that Quokka can effectively leverage LLMs to generate invariants that meaningfully accelerate verification.
-The consistent speedup gains across different model families, from 7B parameter models to frontier models, and the progressive improvement with model scale underscore the effectiveness of our approach.
-Correctness vs. Performance Gap.
-A striking observation is the substantial gap between generating correct invariants and achieving meaningful speedups.
-For example, gpt-5.2 produces 710 correct invariants but achieves $2.0\times$ speedup on only 51 instances, while claude-opus-4.5 generates 689 correct invariants yet delivers $2.0\times$ speedup on merely 34 instances.
-This gap highlights that correctness alone is insufficient: invariants must also be strong enough to meaningfully accelerate verification, requiring not only logical correctness but also strategic insight into which invariants will most effectively guide the solver.
-Model Inference Time.
-We include model inference time in our evaluation to provide a fair and realistic assessment of end-to-end performance.
-As documented in Appendix A.4, smaller models such as Llama-3.1-8B and Qwen2.5-7B exhibit negligible generation times (0.3–0.4 seconds), while most models complete inference within 3 seconds.
-The o3 model has the highest overhead at 8.1 seconds on average.
-Importantly, unlike LOOPY (Kamath et al., 2023), which invokes the LLM up to 15 times per program and thus spends a substantial portion of its time budget on API calls, our approach requires only a single LLM invocation per program.
-Consequently, LLM inference accounts for only a small fraction of total runtime.
-Furthermore, our evaluation dataset contains sufficiently challenging programs (with solving times up to 600 seconds) such that solver time naturally dominates inference overhead, making the comparison meaningful even when inference latency is included.
-Prompting-Based Methods.
-We investigate the impact of different prompting strategies on invariant generation quality using gpt-5.2.
-As detailed in Appendix A.6, we evaluate 0-shot, 2-shot, 4-shot, and chain-of-thought (CoT) prompting.
-Few-shot prompting does not consistently improve performance, while CoT prompting yields the weakest results, likely due to increased generation time without proportional quality gains.
-These findings indicate that simple zero-shot prompting is most effective for invariant synthesis.
-5.2 Comparison with Prior LLM-based Verifiers
-Method # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
-LaM4Inv 1 1 1 1 1
-LOOPY 8 7 5 3 1
-LEMUR-original 19 16 13 12 10
-LEMUR-gpt-5.2 39 30 28 25 22
-Quokka-gpt-5.2 81 65 60 56 51
-Table 3: Comparison of speedup performance across different methods.
-We report the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over UAutomizer.
-Table˜3 compares the speedup performance of Quokka against prior LLM-based verifiers across different speedup thresholds, while Figure˜2 shows the number of solved instances over time for each method.
-Quokka Outperforms Prior LLM-based Verifiers.
-Quokka with gpt-5.2 achieves state-of-the-art results across all speedup metrics.
-Among prior work, LEMUR (Wu et al., 2024b) represents the strongest baseline.
-To ensure a fair comparison, we enhance LEMUR by replacing its original model (gpt-4) with gpt-5.2, creating LEMUR-gpt-5.2.
-Even with this enhancement, Quokka-gpt-5.2 significantly outperforms LEMUR-gpt-5.2: at the $1.2\times$ speedup threshold, Quokka-gpt-5.2 accelerates 81 instances versus 39 for LEMUR-gpt-5.2 and 19 for LEMUR-original.
-At $2.0\times$ speedup, Quokka-gpt-5.2 achieves 51 instances compared to 22 for LEMUR-gpt-5.2 and 10 for LEMUR-original.
-LaM4Inv (Wu et al., 2024a) and LOOPY (Kamath et al., 2023) demonstrate minimal speedup capabilities, achieving at most 1 and 8 instances respectively at the $1.2\times$ level.
-These results demonstrate that Quokka establishes a new state-of-the-art for LLM-based verification acceleration.
-Differences in Base Solvers.
-Different LLM-based verification frameworks are built on different base solvers: Quokka and LEMUR use UAutomizer, LOOPY uses Frama-C (Cuoq et al., 2012), and LaM4Inv uses ESBMC (Gadelha et al., 2018).
-The choice of base solver significantly impacts the performance, as stronger solvers provide a better foundation for speedup.
-Since UAutomizer represents the current state-of-the-art solver, we believe future work should prioritize developing LLM-based verifiers atop state-of-the-art solvers such as UAutomizer.
-For a direct comparison between UAutomizer and prior LLM-based solvers, we also refer readers to Appendix A.3, where we evaluate UAutomizer on the dataset released by each of the prior works.
-Algorithmic Simplicity and Parallelism.
-Prior work predominantly adopts complex, sequential, and iterative processes with highly customized algorithms to handle LLM-generated invariants.
-In contrast, Quokka employs a simple, verifier-based, and parallel approach that directly leverages the base solver’s capabilities.
-LaM4Inv uses a “query-filter-reassemble” strategy that filters and combines model-generated predicates through a complex algorithm, rather than using them directly.
-LOOPY employs a highly-customized Houdini algorithm with iterative fixing to filter candidates.
-LEMUR is closest to our approach, also directly using the base solver to verify the correctness of invariants, but follows a purely sequential decision process with backtracking.
-In contrast, Quokka directly uses LLM-generated invariants without decomposition or reassembly, and exploits parallelism by issuing the verification queries in parallel.
-This design reduces implementation complexity while enabling more efficient hardware utilization, contributing to the superior speedup performance demonstrated in our experiments.
-Figure 2: Number of instances solved by different methods over varying timeout budgets.
-Model # Incorrect Invariants # Goal Timeout # Invariant Timeout # Both Timeout
-gpt-5.1 78 96 245 110
-claude-sonnet-4 110 160 209 145
-gpt-5.2 56 139 230 121
-Table 4: Failure mode breakdown for the top three models on Quokka.
-Incorrect Invariants: the invariant is refuted by the solver.
-Goal Timeout: the invariant is verified, but proving the assertion under it times out.
-Invariant Timeout: verifying the invariant itself times out.
-Both Timeout: both verification queries time out.
-Model # Correct Invariants # Fast1.0 # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
-Qwen2.5-7B (base) 419 116 34 26 23 20 18
-Qwen2.5-7B (finetuned) 431 141 39 27 26 24 20
-Table 5: Performance comparison of base and fine-tuned Qwen2.5-7B. Fine-tuning leads to modest improvements in both correctness and speedup metrics.
-5.3 Failure Mode Analysis
-To understand why many invariants fail to accelerate verification in Quokka, we analyze failure modes for the top three models.
-We categorize failures into four types: 1) Incorrect Invariants: the invariant is refuted by the solver; 2) Goal Timeout: the invariant is verified, but proving the assertion under it times out; 3) Invariant Timeout: verifying the invariant itself times out; 4) Both Timeout: both verification queries time out.
-The results in Table˜4 reveal that while some failures stem from incorrect invariants, the dominant issue is timeout-related.
-This suggests that current models lack an understanding of what makes verification queries tractable for symbolic solvers.
-Future work should explore training reward models or other mechanisms to help models internalize solver difficulty, enabling them to propose invariants that genuinely simplify verification by decomposing hard problems into easier subproblems rather than inadvertently generating verification conditions that are as difficult or harder to solve than the original proof goal.
-Figure 3: Effect of Best-of-N sampling on the number of instances under different Fastp settings.
-5.4 Results of Fine-Tuning and Best-of-N Sampling
-Modest Improvement from Supervised Fine-Tuning.
-We perform supervised fine-tuning using LoRA (Hu et al., 2022) on Qwen2.5-7B for 3 epochs.
-As shown in Table˜5, fine-tuning yields modest improvements: correct invariants increase from 419 to 431, and instances achieving meaningful speedups improve across all thresholds.
-For example, Fast1.0 improves from 116 to 141, and Fast1.2 improves from 34 to 39.
-These results suggest that domain-specific fine-tuning can improve both correctness and speedup, though gains remain moderate.
-Benfits of Best-of-N Sampling.
-We evaluate Best-of-N sampling by generating multiple invariant candidates from gpt-5.2 at temperature 0.7 and selecting the one yielding the fastest verification time.
-As shown in Figure˜3, increasing N from 1 to 4 consistently improves speedup across all Fastp thresholds.
-However, performance plateaus beyond N=8 due to hardware constraints: our 20-core CPU must verify 2N queries in parallel, and when N exceeds 10, resource contention may negate sampling benefits.
-Memory limits can further constrain scalability.
-While Best-of-N sampling is effective, its benefits are hardware-bounded, though modest values of N yield meaningful improvements.
-6 Conclusion
-This work introduced Quokka, a simple yet effective framework for LLM-based invariant synthesis that accelerates program verification.
-Our approach employs a principled verification procedure that decomposes the original proof goal into two independent subproblems and launches verifier queries in parallel, providing sound evaluation while achieving state-of-the-art speedup results.
-Using a benchmark of 866 instances, we evaluated 9 state-of-the-art LLMs across multiple model families and compared against prior LLM-based verifiers.
-The results demonstrate that Quokka consistently outperforms all existing LLM-based verifiers, achieving speedups of at least 1.2× on 81 instances compared to 39 instances for the previous best approach.
-We further demonstrated that supervised fine-tuning and Best-of-N sampling can yield measurable improvements in accelerating verification.
-Our work establishes a new state-of-the-art for LLM-based program verification and provides a principled foundation for future research in this direction.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A Appendix
-A.1 Prompt Template Used for Training Dataset Generation
-Table˜A1 shows the prompt template used for synthesizing training programs from seed programs.
+A.3 Prompt Template Used for Training Dataset Generation
+Table A1 shows the prompt template used for synthesizing training programs from seed programs.
 You will be shown 3 example C programs.
@@
 Table A1: Prompt template for synthetic data generation.
-A.2 An Example from the Fine-Tuning Dataset
+A.4 Prompt Template Used for Invariant Generation
+Table A2 shows the prompt template used to prompt the LLM to generate a loop invariant for a given C program.
+System prompt:
+Given a C program with line numbers, you need to identify the loop invariant to be added at one of the provided points in the program.
+A loop invariant is an expression that is true before, during, and after the execution of a loop.
+You will be provided the set of points in the program where loop invariants can be added. Each point will be specified as after line X.
+Choose one of the points to insert a loop invariant. Respond with exactly this format:
+"After line X, insert assume(condition);"
+Where:
+- X refers to the provided line number where the loop invariant needs to be added
+- condition is the loop invariant condition you want to insert. The loop invariant should not modify any variable (no ++, --, +=, -=, =).
+- only output one loop invariant
+User prompt:
+Program:
+{PROGRAM}
+Choose one of the points to insert a correct loop invariant at in the program:
+{POINTS}
+Please analyze the provided C program and output only where to insert a loop invariant in the specified format.
+Table A2: Prompt template used to generate a loop invariant for a given program.
+The placeholder {PROGRAM} denotes the input C program with line numbers, and {POINTS} denotes the candidate locations where the invariant may be inserted.
+A.5 An Example from the Fine-Tuning Dataset
 Original Program (Input to UAutomizer)
@@
 Figure A1: An example from the fine-tuning dataset: program and its loop invariant generated by UAutomizer.
-Figure˜A1 shows an example from the fine-tuning dataset with the program to the UAutomizer and the generated loop invariant.
+Figure A1 shows an example from the fine-tuning dataset with the program to the UAutomizer and the generated loop invariant.
 The loop invariant holds at Line 6 (the beginning of the loop).
@@
 Thus, the invariant $I$ indeed suffices to prove the final assertion.
-A.3 Results of UAutomizer on Prior Datasets
+A.6 Results of UAutomizer on Prior Datasets
 Verifier Total Instances Solved Instances under Different Timeouts
@@
 Nevertheless, for our work, we focus on problems that UAutomizer can solve within the time budget, as our primary objective is to quantify speedup.
-A.4 Inference Time of Different Models
-Table˜A3 presents the average inference time per instance for different LLMs when generating invariants.
+A.7 Inference Time of Different Models
+Table A4 presents the average inference time per instance for different LLMs when generating invariants.
 As expected, smaller open-source models such as Llama-3.1-8B and Qwen2.5-7B exhibit the fastest generation times, averaging around 0.3–0.4 seconds per instance.
@@
 Table A4: Average inference time per instance for different LLMs when generating invariants.
-Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
-gpt-5.1 399 10 9 9 9 9
-gpt-5.2 455 17 11 10 9 9
+Model # Correct Invariants 60s 500s
+$\Delta$ # Solved $\overline{T}$ $\Delta$ # Solved $\overline{T}$
+gpt-5.1 399 31 394 34.8 1 399 272.9
+gpt-5.2 455 30 450 31.2 1 455 240.9
 Table A5: Performance of different LLMs when using ESBMC as the base verifier.
-We report the number of instances with verified-correct invariants and the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over ESBMC.
-A.5 Results for ESBMC as the Base Solver
+A.8 Results for ESBMC as the Base Solver
 To demonstrate the generality of Quokka’s methodology, we conduct additional experiments using ESBMC (Gadelha et al., 2018) as the base solver instead of UAutomizer.
-As shown in Table˜A4, Quokka successfully accelerates ESBMC as well, with both gpt-5.1 and gpt-5.2 generating hundreds of correct invariants and achieving speedups on multiple instances.
-For example, gpt-5.2 produces 455 correct invariants and achieves speedups of at least $1.2\times$ on 17 instances.
+As shown in Table A5, Quokka successfully accelerates ESBMC as well, with both gpt-5.1 and gpt-5.2 generating hundreds of correct invariants and solving additional instances.
 These results confirm that Quokka is not limited to a specific solver and can be applied to accelerate different verification tools.
-A.6 Results for Prompting-Based Methods
-Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
-gpt-5.2 (CoT) 639 41 32 26 22 20
-gpt-5.2 (4-shot) 756 72 62 53 47 43
-gpt-5.2 (2-shot) 764 83 67 58 53 49
-gpt-5.2 710 81 65 60 56 51
+A.9 Results for Prompting-Based Methods
+Model # Correct Invariants 60s 500s
+# $\Delta$ # Solved $\overline{T}$ # $\Delta$ # Solved $\overline{T}$
+gpt-5.2 (CoT) 639 18 609 33.4 2 638 153.5
+gpt-5.2 710 21 681 22.2 1 710 105.1
+gpt-5.2 (4-shot) 756 20 721 20.1 0 754 81.1
+gpt-5.2 (2-shot) 764 20 730 19.4 2 763 76.0
 Table A6: Comparison of different prompting strategies for gpt-5.2.
 We investigate the impact of different prompting strategies on invariant generation quality using gpt-5.2.
-As shown in Table˜A5, we compare chain-of-thought (CoT) prompting, few-shot prompting with 2 and 4 examples, and our default zero-shot approach.
-The results show that few-shot prompting does not necessarily improve performance and may introduce noise or distract the model from the task at hand.
-Chain-of-thought prompting yields the weakest results across all metrics, with only 639 correct invariants and 41 instances achieving $1.2\times$ speedup.
+As shown in Table A6, we compare chain-of-thought (CoT) prompting, few-shot prompting with 2 and 4 examples, and our default zero-shot approach.
+The results show that few-shot prompting can sometimes improve performance (e.g., 2-shot is generally better than 0-shot), but more examples may also introduce noise or distract the model from the task at hand (e.g., 4-shot is worse than 2-shot).
+Chain-of-thought prompting yields the weakest results across all metrics, with only 639 correct invariants and the fewest solved instances.
 This suggests that explicit reasoning steps may increase generation time without proportional benefits to invariant quality, thereby reducing overall speedup.
-These findings indicate that prompting-based strategy can hardly improve the task of invariant synthesis.
+These findings indicate that few-shot prompting could improve the task of invariant synthesis.
+A.10 Detailed Comparison against Prior Work
+Comparison of Base Solver
+Different LLM-based verification frameworks are built on different base solvers: Quokka and LEMUR use UAutomizer, LOOPY uses Frama-C (Cuoq et al., 2012), and LaM4Inv uses ESBMC (Gadelha et al., 2018).
+The choice of base solver significantly impacts the performance, as stronger solvers provide a better foundation for speedup.
+Since UAutomizer represents the current state-of-the-art solver, we believe future work should prioritize developing LLM-based verifiers atop state-of-the-art solvers such as UAutomizer.
+For a direct comparison between UAutomizer and prior LLM-based solvers, we also refer readers to Appendix A.6, where we evaluate UAutomizer on the dataset released by prior works.
+Detailed Comparison with LEMUR.
+To better understand the performance differences between LEMUR and Quokka, we analyze their implementation strategies and examine the instances where Quokka outperforms LEMUR.
+LEMUR allows LLMs to insert invariants at multiple locations (both before and at the beginning of loops) and generates multiple candidate invariants for each location.
+It then sequentially evaluates each candidate, verifying both the original assertion and the invariant itself.
+If verification times out within 30 seconds, LEMUR recursively invokes the LLM again to generate additional invariants.
+In contrast, Quokka adopts a much simpler strategy: it restricts invariant placement to the beginning of loops, asks the LLM to select a single location (when multiple loops exist), and generates only one invariant.
+The two verification queries (for the assertion and the invariant) are then executed in parallel without recursion or additional LLM invocations.
+Among the 353 instances that Quokka solves but LEMUR does not, 129 instances (36.5%) contain multiple loops.
+We hypothesize that for programs with multiple loops, LEMUR’s exhaustive search strategy becomes less effective due to the combinatorially larger space of invariant placements and candidates it must explore.
+In contrast, Quokka’s focused approach (i.e., requiring the LLM to commit to a single location and invariant) appears more effective in these complex scenarios.
+We identify three key advantages of our design.
+First, parallel execution of verification queries reduces wall-clock time by simultaneously checking both the assertion and the invariant.
+Second, constraining the LLM to select a single location and invariant forces more deliberate reasoning about which invariant will be most impactful, rather than relying on exhaustive search over many candidates.
+Third, eliminating recursion and repair loops reduces both implementation complexity and the risk of execution failures (which we observe in LEMUR’s logs).
+Together, these design choices enable Quokka to achieve superior performance while maintaining a simpler and more robust implementation.
+Model # Incorrect # Goal Timeout # Invariant Timeout # Both Timeout
+gpt-5.1 78 96 245 110
+claude-sonnet-4 110 160 209 145
+gpt-5.2 56 139 230 121
+Table A7: Failure mode breakdown for the top three models on Quokka.
+Incorrect: the invariant is refuted by the solver.
+Goal Timeout: the invariant is verified, but proving the assertion under it times out.
+Invariant Timeout: verifying the invariant itself times out.
+Both Timeout: both verification queries time out.
+A.11 Failure Mode Analysis
+As shown in Table A7, we categorize failures into four types: 1) Incorrect Invariants: the invariant is refuted by the solver; 2) Goal Timeout: the invariant is verified, but proving the assertion under it times out; 3) Invariant Timeout: verifying the invariant itself times out; 4) Both Timeout: both verification queries time out.
+A.12 Case Studies: Verification Timeout in Quokka
+To better understand the verification timeouts in Quokka, we manually examined 50 timeout cases and identified five recurring categories.
+In this section, we select two of these categories for closer analysis and present one representative program from each.
+The first category includes cases where the LLM proposes a trivial but correct invariant, yet the solver still cannot complete verification within the timeout budget.
+The second category includes cases where the LLM proposes a correct invariant that is almost identical to the original assertion.
+Case 1: Trivial and correct invariant.
+Figure A2 shows a benchmark where the LLM proposes the invariant c == y after line 12.
+This invariant is correct because c and y are both initialized to zero before the loop and are incremented by exactly one in every iteration, so the equality is preserved throughout the loop.
+Moreover, it is trivial to prove: its validity follows directly from the initialization and the loop updates, and can be established by a simple inductive argument.
+Nevertheless, in this case, the solver still cannot prove this invariant within the timeout budget.
+This example shows that verification timeout may occur even when the proposed invariant is both semantically correct and trivial to prove.
+1int main() {
+2 short k;
+3 long long y, x, c;
+4 k = __VERIFIER_nondet_short();
+5 assume_abort_if_not(k >= 0 && k <= 5);
+6 assume_abort_if_not(k <= 256);
+7
+8 y = 0;
+9 x = 0;
+10 c = 0;
+11
+12 while (1) {
+13 __VERIFIER_assert(-2 * y * y * y * y * y * y - 6 * y * y * y * y * y - 5 * y * y * y * y + y * y + 12 * x == 0);
+14
+15 if (!(c < k)) {
+16 break;
+17 }
+18 c = c + 1;
+19 y = y + 1;
+20 x = y * y * y * y * y + x;
+21 }
+22
+23 return 0;
+24}
+Figure A2: A timeout case where the proposed invariant c == y is semantically correct and trivial to prove, yet the solver still cannot establish it within the timeout budget.
+Case 2: Correct invariant as a trivial variant of the original assertion.
+A different phenomenon is shown in Figure A3, where the LLM proposes the invariant
+$x=\frac{y(y+1)}{2}.$
+This invariant is correct: it holds initially and is preserved by the loop updates. However, it is essentially just a trivial algebraic reformulation of the original assertion
+$y^{2}-2x+y=0.$
+This possibly suggests a limitation of the LLM in decomposing the verification task into a genuinely simpler intermediate property, instead of merely restating the target assertion in an equivalent form.
+1int main() {
+2 int k;
+3 long long y, x, c;
+4 k = __VERIFIER_nondet_int();
+5 assume_abort_if_not(k >= 0 && k <= 10);
+6
+7 y = 0;
+8 x = 0;
+9 c = 0;
+10
+11 while (1) {
+12
+13 if (!(c < k)) {
+14 break;
+15 }
+16
+17 c = c + 1;
+18 y = y + 1;
+19 x = y + x;
+20 }
+21 __VERIFIER_assert((y * y) - 2 * x + y == 0);
+22 return 0;
+23}
+Figure A3: A case where the proposed invariant is correct but is essentially a trivial algebraic variant of the original assertion.